### v1

# Visual Document Understanding and Question Answering: A Multi-Agent Collaboration Framework with Test-Time Scaling

Abstract
Existing vision-language models (VLMs), whether generalists or specialists, remain constrained by their parameter scale, lack robust self-correction capabilities, and underperform in tasks involving long visual contexts and complex reasoning, resulting in suboptimal performance on document-based tasks. To address this, we propose MACT, a Multi-Agent Collaboration framework with Test-Time scaling, tailored for visual document understanding and visual question answering (VQA). It comprises four distinct small-scale agents, i.e., planning, execution, judgment, and answer agents, with clearly defined roles and effective collaboration. Notably, the judgment agent exclusively verifies correctness and redirects to prior agents for revisions, outperforming conventional correction strategies. To further expand the capability boundaries of the framework, we propose mixed reward modeling that balances agent-specific abilities and global collaboration, as well as agent-wise hybrid test-time scaling, which customizes different scaling strategies for each agent based on their functions. Evaluated on benchmarks spanning both document-based and non-document-based settings, our MACT shows superior performance with a smaller parameter scale without sacrificing the ability of general and mathematical tasks. Especially, it stands out in benchmarks involving long visual contexts and complicated reasoning. The three variants of MACT consistently hold the top three positions in average scores, leading in 13 of the 15 benchmarks. Code will be available at: https://github.com/YU-deep/MACT.git.
Introduction
For humans, acquiring and understanding visual information from documents—such as texts, webpages, charts, and tables—is a common and indispensable part of daily life. As a derivative task, document-based VQA provides an assessment of the understanding of both textual and visual document information as well as their relations, whose performance is quantifiable. Therefore, numerous studies aim to propose benchmarks for more complex document-based tasks of diverse types to evaluate model performance fairly and comprehensively. Specifically, the types have expanded from single-page documents, simple charts, and tables to real-world multi-page documents, complicated webpages, bar charts, pie charts, line charts, composite tables, matrix tables, and mathematical reasoning. Additionally, the form of the questions has also transformed from straightforward Yes/No, single-choice or direct information extraction questions to diversified question types, including but not limited to visual details detection, cross-page or multi-hop information extraction and summarization, comparison, multi-step calculation, intersection finding and unanswerable questions, which are answered based on the layout, structure, visual and textual contents. Thus, the latest flagship VLMs of various series are dedicated to expanding these boundaries of understanding documents.
(a) Qwen2.5-VL Series Based
(b) MiMo-VL Series Based
(c) InternVL-3 Series Based
Figure 1: Comparisons among the three variants of our MACT, their base models, and larger models in the same series.
Recently, the latest released general VLMs have showcased exceptional understanding and generalization capabilities. For instance, closed-source models, such as GPT-4o (Hurst et al. 2024), Gemini-2.0-Pro (DeepMind 2025) and Claude-3.7-Sonnet (Anthropic 2024) have surpassed human-level scores in certain tasks, while open-source models, e.g., Llama-3.2-Vision (Grattafiori et al. 2024), Qwen2.5-VL (Bai et al. 2025), MiMo-VL (Xia et al. 2025), InternVL-3 (Zhu et al. 2025) and Ovis2 (Lu et al. 2024b) have also achieved great performance. However, their performances in document-based tasks remain suboptimal due to the biases or deficiencies in the training data. To further enhance the ability, some researchers have endeavored to develop specialized VLMs. For example, UReader (Ye et al. 2023), TextMonkey (Liu et al. 2024b) and mPLUG-DocOwl2 (Hu et al. 2024b) focus on document understanding. However, their overall performance lags far behind that of the latest general VLMs with larger parameters.
To summarize, currently existing VLMs, including both generalists and specialists, exhibit three key limitations in understanding visual documents and answering questions: (1) Constrained by Parameter Scale: There are substantial performance gaps between larger-parameter VLMs and smaller counterparts in the document domain, and the potential of smaller-parameter VLMs remains underactivated. (2) Deficiency in Self-Correction Capacity: As self-correction is indispensable when humans tackle tasks involving challenging and complex documents, existing methods either fail to adequately address this capability or adopt suboptimal designs. (3) Underperformance on Long Visual Contexts and Reasoning: Results on some document-based benchmarks, which feature long visual contexts or require dense reasoning, are unsatisfactory, with notably low accuracy.
Given these issues, we propose a multi-agent collaboration framework with a dedicated test-time scaling strategy for visual document understanding and question answering. Within this framework, we employ four relatively lightweight collaborative agents: planning, execution, judgment, and answer agents. Through their collaboration and mixed reward modeling, each agent directs attention to both agent-tailored and global objectives, thereby unleashing the potential of the agents with small-scale parameters. For more efficient self-correction capacity, we design a mechanism that employs an additional judgment agent. It only checks the correctness of the generated execution plans and execution processes, with corrections implemented by the preceding agents. Additionally, we introduce an agent-wise hybrid test-time scaling approach that considers the distinct functions and features of each agent, markedly improving abilities in handling long visual contexts and complex reasoning. We train three variants of MACT based on different groups of base models using a two-stage pipeline, and evaluate them on 15 benchmarks, comprising 10 document-based and 5 non-document-based ones, to comprehensively assess document understanding as well as general and reasoning performance. Compared to models of comparable scale and larger-scale ones, the results indicate that the three variants of MACT show an average increase of at least 3.2% to 5.6%, securing the top three positions and achieving the best performance in 13 out of 15 benchmarks. Furthermore, our models have distinct advantages in long visual contexts and reasoning tasks, as further validated by results on selected non-document benchmarks.
Methodology
This section provides a detailed elaboration of our proposed MACT, divided into three subsections. We first formally define the task, then introduce the four agent components: the planning agent, the execution agent, the judgment agent, and the answer agent, as well as the collaboration mechanisms among them. Finally, we detail the mixed reward modeling and agent-wise hybrid test-time scaling tailored to our multi-agent collaboration framework.
Task Definition
For the question answering task defined in this work, given a question $\mathcal{Q}$ and the its corresponding visual inputs $\mathcal{V}=\{v_{1},v_{2},\dots,v_{n}\}$, the target is to generate a correct answer output $\mathcal{O}$. This answer should leverage the information within $\mathcal{V}$ and answer the proposed question comprehensively and accurately.
Figure 2: The overview of MACT. The upper part demonstrates our proposed multi-agent framework, which has four collaborative agents for tailored tasks. When the judgment agent detects mistakes, it redirects to previous agents for corrections. While the lower part illustrates the mixed reward modeling and agent-wise hybrid scaling for the multi-agent framework.
Multi-Agent Collaboration Framework
As shown in the upper part of Figure 2, MACT consists of four collaborative agents: sequentially planning agent ($\mathcal{A}_{plan}$), execution agent ($\mathcal{A}_{exe}$), judgment agent ($\mathcal{A}_{judg}$), and answer agent ($\mathcal{A}_{ans}$). Each agent is role-specialized, which accomplishes its functionality and then passes the generated contents to the subsequent agent or outputs the final answers.
[Uncaptioned image] Planning Agent.
$\mathcal{A}_{plan}$ mainly focuses on the analysis and decomposition of the original question, making high-level executive plans. Largely following the analogical prompting principles (Yasunaga et al. 2024), it first generates $\mathit{N_{p}}$ relevant sample problems along with their corresponding plans simultaneously. By referring to each sample problem and its corresponding plan, it produces distinct execution plans for the current problem respectively, providing diverse pathways to accomplishment. It should be noted that this agent only yields high-level execution plans without implementation details. It ensures that $\mathcal{A}_{plan}$ can better formulate the plan from a holistic perspective, and avoid interfering with the execution of $\mathcal{A}_{exe}$. Moreover, it avoids premature binding to specific tools, enabling $\mathcal{A}_{exe}$ to select optimal resources for diverse tasks dynamically.
[Uncaptioned image] Execution Agent.
$\mathcal{A}_{exe}$ is designed to execute plans step by step and output the execution process using the selected tools from the tool library. Specifically, it breaks down the steps in the plan into execution units, with information for each unit populated into a template. $\mathcal{A}_{exe}$ then executes these units sequentially, deriving the intended outputs from each. Once all steps are completed, it concatenates the full output and the execution processes of all steps in sequence, passing them to subsequent agents.
[Uncaptioned image] Judgment Agent.
To enhance the quality of completing challenging tasks, humans often judge and correct mistakes that arise in planning or execution—such as reversed sequences, incorrect dependencies, redundant or missing steps, mismatches between plans and processes, and disconnects between steps. Similarly, some multi-agent systems, primarily focused on code generation, incorporate mechanisms that either (a) internally correct errors within the same agent (Yang et al. 2024; Sun et al. 2024) or (b) deploy an additional agent to handle both judgment and correction (Islam, Ali, and Parvez 2024; Li et al. 2025a). However, the internal correction mechanism (a) struggles to identify most mistakes and may fall into cognitive blind spots, as generation and correction rely on the same model. In contrast, the latter approach (b) requires the agent to possess strong capabilities for both judgment and correction, thereby necessitating models with larger parameters and more complex reward modeling designs. Furthermore, using a separate agent to regenerate mistaken components may result in incompatibilities or even conflicts with existing parts, given the autonomy of the two agents. In RL, the optimization objective for both (a) and (b), as illustrated in Figure 3, is to pass verification. This can lead to strategic production of vague statements or the omission of details, resulting in corrections that appear correct superficially but are actually misleading.
To address the limitations of both approaches, we design a novel judgment strategy for multi-agent systems, which introduces an additional judgment agent to separate judgment from correction, thereby fostering specialization of labor. More specifically, $\mathcal{A}_{judg}$ is primarily responsible for assessing the correctness of steps in previously generated execution plans and processes, without engaging in direct correction. If mistakes are detected in any step, $\mathcal{A}_{judg}$ identifies the specific problematic step, provides a brief description of the error, and routes the issue to the appropriate agent—either $\mathcal{A}_{plan}$ or $\mathcal{A}_{exe}$—for correction. For each plan-process pair that is already correct or has been corrected, $\mathcal{A}_{judg}$ forwards them to $\mathcal{A}_{ans}$. This strategy, by decoupling judgment from correction, introduces a neutral judge with reduced subjective bias, whose focus remains solely on judgment rather than correction. Additionally, the reward design for the judgment agent can be intuitively simplified, without evaluating correction outcomes. To prevent infinite correction loops, the maximum number of corrections $N_{c}$ is set to 3.
Figure 3: Comparisons of (a) internal correction, (b) an extra agent for both judgment and correction, and (c) our strategy utilizing an independent judgment agent.
[Uncaptioned image] Answer Agent.
Counterintuitively, $\mathcal{A}_{ans}$ incorporates both the correct execution process and incorrect segments from prior processes. This facilitates direct focus on modifications within the corrected process, thereby preventing the omission of error-prone details.
Collaboration.
Throughout the full implementation of the multi-agent collaboration framework, visual inputs and questions are first fed into $\mathcal{A}_{plan}$, whose generated plan is then executed by $\mathcal{A}_{exe}$. $\mathcal{A}_{judg}$ judges the correctness the resulting execution plan-process pairs and outputs mistake flags. For correct pairs, $\mathcal{A}_{ans}$ outputs the final answers, while the information of mistakes will be passed to the previous agent $\mathcal{A}_{plan}$ or $\mathcal{A}_{exe}$ for correction, and the procedure repeats. For clarity in illustrating the workflow, we set the number of generated relevant plans to 1 and simplify the test-time scaling designs in Algorithm 1, and these details are elaborated in the subsequent section.
Algorithm 1 Multi-Agent Collaboration Procedure
0: question $\mathcal{Q}$, visual inputs $\mathcal{V}=\{v_{1},v_{2},\dots,v_{n}\}$, planning agent $\mathcal{A}_{plan}$, execution agent $\mathcal{A}_{exe}$, the tool library of the former agent $\mathcal{T}=\{t_{1},t_{2},\dots,t_{n}\}$, judgment agent $\mathcal{A}_{judg}$, answer agent $\mathcal{A}_{ans}$, maximum number of corrections $N_{c}$
0: answer output to the question $\mathcal{O}$
1: Initialize the four agents: $\mathcal{A}_{plan}$, $\mathcal{A}_{exe}$, $\mathcal{A}_{judg}$, $\mathcal{A}_{ans}$
2: Initialize $N_{c}=3,t=0$
3: Initialize prompts formatter $\mathcal{PF}$
4: Initialize $flag_{plan}=flag_{exe}=false,\mathcal{M}=empty$
5:while true do
6:  $p_{1}\leftarrow\mathcal{PF\left(\mathcal{Q}\right)}$
[…]
8:  $p_{2}\leftarrow\mathcal{PF}\left(\mathcal{Q},\mathcal{P}_{relevant},\mathcal{M}\right)$
[…]
10:  for each step $s_{i}$ in $\mathcal{P}$ do
11:   $p_{3}\leftarrow\mathcal{PF}\left(\mathcal{Q},s_{i},\mathcal{T},\mathcal{M}\right)$
[…]
13:  end for
[…]
15:  $p_{4}\leftarrow\mathcal{PF\left(\mathcal{Q},\mathcal{P},\mathcal{E}\right)}$
[…]
18:  if not $flag_{plan}$ and not $flag_{exe}$ or $t\geq N_{c}$ then
19:   break
20:  else if $flag_{plan}$ then
21:   $t\leftarrow t+1$
[…]
23:  else if $flag_{exe}$ then
24:   $t\leftarrow t+1$
[…]
26:  end if
27:end while
[…]
Mixed Reward modeling
To guide RL in our multi-agent system and further enhance both the specialized and collaborative capabilities of each agent, we design a mixed reward strategy that integrates agent-specific and global outcome reward signals. Given that the four agents in the framework have distinct functions and reward signal preferences, we first incorporate agent-tailored rewards for each. In particular, for $\mathcal{A}_{plan}$ and $\mathcal{A}_{exe}$, whose outputs are step-by-step, we employ an effective multi-modal reward model to yield step-wise process reward signals. These step-wise rewards provide instant feedback for each step and hierarchical rewards for the two agents, boosting accuracy on challenging tasks. For $\mathcal{A}_{judg}$ and $\mathcal{A}_{ans}$, we directly use a reward model to generate one reward signal for each output of each agent. Furthermore, we apply a global reward, computed based on the final selected answer along the chosen path of the four agents. It strengthens rewards for correct paths, guiding the model to explore valid paths while mitigating adverse effects from erroneous ones. Besides, it mitigates the selfishness of agents to some extent, as individual agents tend to maximize their own rewards without regard for the global objective, and this global reward helps to alleviate such behavior.
Agent-Wise Hybrid Test-Time Scaling
Since that scaling test-time computation can activate the long-context understanding and reasoning capabilities of small-parameter VLMs, dynamically allocating additional computing resources and enabling multiple paths in the test time, we propose an agent-wise hybrid scaling strategy during inference. Existing test-time scaling methods fall into four main categories: parallel scaling, sequential scaling, hybrid scaling, and internal scaling. However, these strategies are originally designed for single models or agents, overlooking the different division of labor among agents and achieving suboptimal performance when applied to multi-agent systems. Accordingly, we propose an agent-wise hybrid scaling strategy tailored to multi-agent architectures, which significantly enhances both agent-specific capabilities and collaborative performance.
For the first three agents with various functions, we customize different test-time scaling strategies to their unique characteristics: (1) Planning Agent: In our design of $\mathcal{A}_{plan}$, it naturally provides relevant sample plans as references for formulating multiple execution plans for the original task. We therefore prompt the generation of $N_{p}$ relevant plans independently, yielding $N_{p}$ parallel paths per query. It establishes a foundation for scaling subsequent agents and increases the likelihood that at least one path will produce the correct answer. (2) Execution Agent: Given that the execution process is divided by step in $\mathcal{A}_{exe}$, we treat each step as an evaluation node. For each node, the agent produces $\mathit{N_{e}}$ candidate executions, which are scored using a pretrained reward model. The top-scoring candidate is selected as the base node for subsequent steps, with all others rejected. (3) Judgment Agent: To judge the correctness of the execution plans and processes, $\mathcal{A}_{judg}$ requires logical analysis and reasoning to comprehensively and accurately detect mistakes. We therefore adopt the budget forcing scaling method (Muennighoff et al. 2025) for this agent, which enforces a minimum number of thinking tokens. When token usage falls below the budget, the agent is encouraged to generate additional thinking tokens, thereby promoting accurate judgments. For the answer agent, whose key function is to summarize prior information and generate the final answer, test-time scaling yields only limited improvements, so we do not apply scaling to it.
Document Non-Document Average
Text Webpage Chart Table General Mathematical
DVQA DUDE SVQA MMLong VisMRC InfVQA ChartQA CharXiv TabVQA TabBen SciQA RealQA MVista MVision MVerse
Average Images Per Question 1.0 5.7 20.0 47.5 1.0 1.2 1.0 1.0 1.0 1.0 0.5 1.0 1.0 1.0 1.0 5.7
Average Question Length 8.3 8.7 9.0 16.4 10.6 11.5 7.2 20.4 10.5 20.3 12.1 11.1 15.6 42.3 35.7 16.0
Average Answer Length 2.1 3.4 2.4 2.8 9.5 1.6 1.2 2.9 4.7 8.5 4.4 1.2 1.2 2.2 1.4 3.3
Generalist
Closed-Source
GPT-4o-latest 93.1 52.7 81.0 40.3 86.4 79.2 86.5 78.7 64.2 51.9 83.3 76.2 62.4 30.7 39.4 67.2
Claude-3.7-Sonnet 94.0 58.1 83.6 33.9 82.5 75.3 92.2 83.5 70.3 54.7 80.4 71.9 69.7 41.2 45.0 69.1
Gemini-2.0-Pro 91.8 54.3 78.9 32.2 91.4 81.6 88.8 83.1 71.2 59.9 80.9 70.5 74.8 54.2 56.6 71.3
Size << 20B
DeepSeek-VL2-4.5B 78.8 - - - 70.4 65.3 74.0 46.9 30.8 - 66.6 60.7 29.5 6.9 7.4 -
LLaVA-OneVision-7B-si 86.6 48.4 52.9 5.3 76.1 74.8 82.3 54.3 47.9 33.3 72.8 62.0 57.9 17.9 19.1 52.8
Qwen2.5-VL-7B-Instruct 94.6 62.3 76.2 22.4 86.8 82.8 87.5 68.2 60.4 49.2 74.0 68.4 67.8 25.5 40.8 64.5
MiMo-VL-7B-SFT 92.9 62.3 74.2 27.5 87.3 82.7 86.1 72.0 57.5 52.5 75.1 67.9 70.7 46.9 53.3 67.3
InternVL3-8B-Instruct 91.1 60.1 72.9 28.6 84.4 78.5 85.4 67.1 59.7 52.4 76.4 66.7 65.5 28.8 38.9 63.8
Llama-3.2-11B-Vision-Instruct 87.3 45.0 66.5 8.8 79.6 73.8 79.6 63.6 50.4 44.6 76.9 60.7 56.5 16.3 32.3 56.1
LlaVa-1.6-vicuna-13B 79.0 42.9 65.6 7.9 77.9 74.7 76.6 52.5 24.9 39.2 70.7 58.1 50.2 11.6 27.5 50.6
InternVL3-14B-Instruct 90.7 61.0 74.5 28.9 87.1 83.1 86.2 75.1 58.2 51.4 77.1 66.4 66.4 36.6 43.5 65.7
Ovis2-16B 92.4 59.6 73.8 29.0 88.5 78.1 87.9 70.7 59.5 46.8 72.5 67.9 69.6 35.6 45.0 65.1
30B $\leq$ Size << 40B
Qwen2.5-VL-32B-Instruct 94.9 66.4 75.0 22.9 86.4 83.8 87.7 75.9 62.7 51.0 74.8 71.7 65.8 32.6 46.4 66.5
LlaVa-1.6-34B 87.8 50.1 69.5 9.9 80.1 78.4 85.3 60.4 32.9 44.4 76.2 67.0 56.7 15.8 17.0 55.4
Ovis2-34B 93.6 63.1 72.4 31.0 85.6 79.1 88.9 71.9 61.6 47.1 72.0 70.7 74.0 31.4 49.5 66.1
InternVL3-38B-Instruct 95.0 61.7 76.3 30.2 89.9 84.6 88.0 79.8 63.1 54.1 77.5 72.8 75.2 34.0 47.6 68.7
70B $\leq$ Size << 100B
LLaVA-OneVision-72B-si 91.7 62.7 63.9 10.6 81.4 75.2 86.1 63.8 52.8 41.5 75.1 70.5 72.5 25.2 27.0 60.0
Qwen2.5-VL-72B-Instruct 95.7 67.0 82.4 24.9 90.3 87.5 89.5 80.6 68.4 58.1 74.4 74.8 77.6 39.8 46.2 70.5
InternVL3-78B-Instruct 95.3 68.9 80.7 32.4 90.7 86.2 89.6 77.9 65.5 57.9 78.0 77.9 79.3 44.3 48.7 71.6
Llama-3.2-90B-Vision-Instruct 91.4 58.6 76.7 19.5 86.0 82.3 86.3 74.4 60.1 49.4 78.5 68.8 58.7 23.4 35.0 63.3
Specialist
MMCA-7B 68.9 - - - 59.0 54.9 64.4 47.5 - - 56.6 52.6 - - - -
UReader-7B 72.3 - - - 70.7 66.6 67.0 52.0 - - 60.4 55.7 - - - -
mPLUG-DocOwl2-8B 87.8 46.5 - - 76.1 72.2 80.8 52.4 25.7 19.8 65.7 60.0 41.3 9.2 14.4 -
TextMonkey-10B 77.2 36.8 - - 74.5 68.1 72.5 52.0 21.7 16.9 65.9 63.4 33.7 5.4 11.7 -
M3DocRAG-10B 79.2 39.5 57.9 29.2 71.8 67.3 73.7 56.5 22.5 16.5 65.4 58.9 41.8 14.8 20.8 47.7
CogAgent-17B 80.9 - - - 70.3 65.3 69.9 54.3 21.7 - 58.8 52.6 - - - -
MDocAgent-39B 86.4 58.4 68.6 36.7 82.8 74.4 82.8 70.2 58.8 48.7 69.8 64.8 46.6 20.1 29.3 59.9
Ours
MACT 96.6 72.5 85.3 43.7 92.0 89.4 91.9 85.2 74.0 57.2 78.5 77.7 81.2 41.8 54.7 74.8
-Qwen2.5-VL-Series-24B +2.0 +10.2 +9.1 +21.3 +5.2 +6.6 +4.4 +17.0 +13.6 +8.0 +4.5 +9.3 +13.4 +16.3 +13.9 +10.3
MACT 94.4 70.8 83.9 47.4 93.8 88.6 91.4 87.2 71.6 62.7 79.2 76.1 85.4 60.1 65.3 77.2
-MiMo-VL-Series-28B +1.5 +8.5 +9.7 +19.9 +6.5 +5.9 +5.3 +15.2 +14.1 +10.2 +4.1 +8.2 +14.7 +13.2 +12.0 +9.9
MACT 96.1 69.8 81.3 46.5 91.0 87.7 90.6 85.0 75.3 61.0 81.3 80.1 83.8 45.8 54.0 75.3
-InternVL3-Series-28B +5.0 +9.7 +8.4 +17.9 +6.6 +9.2 +5.2 +17.9 +15.6 +8.6 +4.9 +13.4 +18.3 +17.0 +15.1 +11.5
Table 1: DVQA, SVQA, MMLong, InfQA, VisMRC, TabVQA, TabBen, SciQA, RealQA, MVista, MVision, and MVerse denote DocVQA, SlideVQA, MMLongBench-Doc, InfographicVQA, VisualMRC, TableVQA-Bench, TableBench, ScienceQA, RealWorldQA, MathVista, Math-Vision, and MathVerse benchmarks, respectively. The values with green, blue, and gray colors denote the best, the second best, and the third best values. Results with - indicate exceeding the maximum context token limit or not being equipped with a specific ability. $\triangle,\Box,\Diamond$ correspond to the base models of MACT for comparisons.
Experimental Settings
Training Pipeline
Since the functions and optimization goals of each agent are independent and diverse, we design a two-stage SFT and RL pipeline for MACT. To understand visual inputs and better accomplish their labors, we choose VLMs for the $\mathcal{A}_{plan}$ and $\mathcal{A}_{exe}$, while LLMs for the $\mathcal{A}_{judg}$ and $\mathcal{A}_{ans}$. For all the four agents, we initiate with pretrained models, and we select three groups of small-parameter base models for different variants of MACT: (1) Qwen2.5-VL Series Based (Bai et al. 2025; Yang et al. 2025): Qwen2.5-VL-7B-Instruct and Qwen2.5-7B/3B-Instruct; (2) MiMo-VL Series Based (Xiaomi 2025; Xia et al. 2025): MiMo-VL-7B-SFT and MiMo-7B-SFT; (3) InternVL3 Series Based (Zhu et al. 2025): InternVL3-9B/8B/2B-Instruct. In the first SFT stage, we initially train a tuned 11B/7B/7B VLM on the selected document-based or non-document-based datasets, mixing data with or without CoT. It aims to enhance their visual understanding and reasoning abilities, providing more robust models for future multi-agent collaboration. Next, we fine-tune an 8B/7B/7B LLM on judgment labels generated via GPT-4o (Hurst et al. 2024) and rule-based verifications. Finally, we fine-tune another 3B/3B/7B LLM on the outputs from preceding agents and ground-truths to enhance its summary ability. The fine-tuned VLM are applied as $\mathcal{A}_{plan}$ and $\mathcal{A}_{exe}$ respectively in the subsequent stage, while the two LLMs function as the $\mathcal{A}_{judg}$ and $\mathcal{A}_{ans}$. In the second RL stage, we generate reward signals based on pretrained reward models and optimize our model via GRPO (Shao et al. 2024). Process reward model VisualPRM (Wang et al. 2025c) is utilized for the step-by-step reward signals of $\mathcal{A}_{plan}$ and $\mathcal{A}_{exe}$, while Skywork-VL-Reward (Wang et al. 2025d) is used for the rewards generation of $\mathcal{A}_{judg}$ and $\mathcal{A}_{ans}$. More training details are in the Appendix.
Datasets
To train our model and comprehensively, objectively assess its visual capabilities, we selected 15 datasets spanning four real-world document types and two non-document types. The four document types are comprised of: (1) Text-Based: DocVQA (Mathew, Karatzas, and Jawahar 2021), DUDE (Van Landeghem et al. 2023), SlideVQA (Tanaka et al. 2023), MMLongBench-Doc (Ma et al. 2025); (2) Webpage-Based: VisualMRC (Tanaka, Nishida, and Yoshida 2021), InfographicVQA (Mathew et al. 2022); (3) Chart-Based: ChartQA (Masry et al. 2022), CharXiv (Wang et al. 2024d); (4) Table-Based: TableVQA-Bench (Kim, Yim, and Song 2024), and TableBench (Wu et al. 2025). To retain generalization and enhance reasoning, two non-document types are involved: (1) General: ScienceQA (Saikh et al. 2022), RealWorldQA (xAI 2024); (2) Mathematical: MathVista (Lu et al. 2024a), Math-Vision (Wang et al. 2024b), and MathVerse (Zhang et al. 2024c). We adhere to the original training and testing splits and use the test splits as evaluation benchmarks. Each instance in these datasets consists of a visual input and a question (except ScienceQA), spanning difficulty levels from easy to hard and encompassing various question types. More details are provided in the Appendix.
Evaluations
Following (Li et al. 2025b), we use the reported initial results of reproducible models where configurations align with ours. For other models, we utilize LMMs-Eval (Zhang et al. 2024b) for fair comparisons on both natively supported and our registered benchmarks. For most benchmarks, we employ GPT-4o (Hurst et al. 2024) as a judge model to evaluate the correctness of the generated answers, which is achieved in LMMs-Eval. For the remaining benchmarks, we utilize their original evaluation metrics, e.g., ANLS and F1. More details can be seen in the Appendix.
Results and Discussions
Main Results
We evaluate MACT across 15 benchmarks: 10 document-based benchmarks to assess its document understanding performance; 2 general benchmarks to gauge generalization ability, mitigating overfitting to specific scenarios; and 3 mathematical benchmarks to evaluate complex reasoning capabilities. For comparison, we select state-of-the-art (SOTA) methods—encompassing both generalist models and document-specific specialists—and categorize them by parameter size.
As shown in Table 1, the MACT-MiMo-VL-Series-28B variant delivers the best average performance, followed by MACT-InternVL3-Series-28B and MACT-Qwen2.5-VL-Series-24B, which rank second and third, respectively. Despite having fewer than 30B parameters, MACT models outperform all comparative methods with under 100B parameters, as well as closed-source models. Notably, MACT-MiMo-VL-Series-28B achieves significant improvements of 5.6% and 5.9% over the top-performing open-source and closed-source models in terms of the average scores. Additionally, the three variants top 13 of the 15 benchmarks, with MACT-MiMo-VL-Series-28B leading on seven. Notably, in MMLongBench-Doc, which features the longest visual context, and across the three mathematical reasoning benchmarks, MACT-MiMo-VL-Series-28B outperforms the second-highest scorer by 7.1%, 10.6%, 5.9%, and 8.7%, respectively. These results highlight the superior overall performance of MACT, as well as its strong capabilities in long-context understanding and reasoning.
Radar charts in Figure 1 compare three variants of MACT with their base models and larger models from the same series. As shown, MACT models significantly outperform their base models by 10.3%, 9.9%, and 11.5% on average across the 15 benchmarks. Compared to their corresponding larger-parameter counterparts, i.e., Qwen2.5-VL-72B-Instruct and InternVL3-78B-Instruct, our MACT-Qwen2.5-VL-Series-24B and MACT-InternVL3-Series-28B achieve average performance gains of at least 6.6% and 3.7%, respectively, with more pronounced improvements observed on long-context and reasoning tasks.
Additional Quantitative Analysis
To further demonstrate the effectiveness of our methodologies, we perform ablation experiments and analyses on the proposed multi-agent collaboration, reward modeling, and test-time scaling strategies. Experiments here are conducted using MACT-Qwen2.5-VL-Series-24B. Specifically, we select three lowest-scoring benchmarks, which are characterized by long visual contexts and reasoning demands, and the average results across all benchmarks to assess their impact on both specific abilities and overall performance.
Ablation Studies
MMLong TabBen MVision Average
w/o Multi-Agent Collaboration 32.5 $\downarrow$11.2 50.8 $\downarrow$  6.4 32.4 $\downarrow$  9.4 66.2 $\downarrow$  8.6
w/o Multi-Agent Collaboration 24.7 $\downarrow$19.0 44.4 $\downarrow$12.8 26.0 $\downarrow$15.8 58.6 $\downarrow$16.2
w/o Mixed Reward Modeling 38.3 $\downarrow$  5.4 54.2 $\downarrow$  3.0 36.7 $\downarrow$  5.1 71.4 $\downarrow$  3.4
w/o Agent-Wise Hybrid Scaling 34.9 $\downarrow$  8.8 50.8 $\downarrow$  6.4 34.5 $\downarrow$  7.3 71.1 $\downarrow$  3.7
MACT 43.7 57.2 41.8 74.8
Table 2: Results of ablations. $\dagger$ denotes a single agent that directly outputs answers, while $\ddagger$ indicates a single agent that incorporates all prompts as a workflow.
$\mathcal{A}_{plan}$ + $\mathcal{A}_{exe}$ $\mathcal{A}_{judg}$ $\mathcal{A}_{ans}$ MMLong TabBen MVision Average
36.3 48.5 34.9 68.4
42.2 56.7 40.2 73.9
36.6 48.6 35.4 68.8
43.7 57.2 41.8 74.8
Table 3: Results of different combinations of agents.
The ablation settings are as follows: (1) w/o Multi-Agent Collaboration: Use a single agent to directly execute tasks and output answers, with our proposed scaling and reward strategies retained. (2) w/o Multi-Agent Collaboration: Use a single agent, which incorporates all prompts from our four proposed agents as a unified workflow, with our proposed scaling and reward strategies retained. (3) w/o Agent-Wise Hybrid Scaling: Retain the multi-agent collaborative framework and apply the mixed reward strategy directly, without test-time scaling. (4) w/o Mixed Reward Modeling: Retain the multi-agent collaborative framework and agent-wise hybrid scaling, without reward signals. As shown in Table 2, the multi-agent collaborative approach contributes the most, outperforming the single-agent system by 8.6%. Notably, consolidating all functions into a single model results in suboptimal performance, which is even worse than the base model, highlighting the necessity of a multi-agent collaborative framework. Both the mixed reward modeling and agent-wise hybrid scaling strategies boost average accuracy by at least 3%. Specifically, the latter yields more significant gains on complex tasks compared to average ones, with especially notable improvements in reasoning tasks.
Analysis of Multi-Agent Collaboration
To assess the utility of each agent and their collaboration, we perform additional experiments on different agent combinations, which are listed in Table 3. The combination of $\mathcal{A}_{plan}$ and $\mathcal{A}_{exe}$ results in a 3.9% average increase over the base model, while adding $\mathcal{A}_{judg}$ further boosts performance by 5.5%. $\mathcal{A}_{ans}$ extraly improves the scores by 0.9%, with this gain expanding to 1.5% for long visual context. Additionally, as in Figure 4(a), we compare our judgment agent strategy with the other two counterparts. Our proposed strategy outperforms the others by at least 2.6% on average, while requiring an average of 0.3 fewer corrections. Moreover, our method achieves optimal correction results when the maximum number of corrections $N_{c}$ is set to 3, whereas the alternatives require $N_{c}$ to be set to 5, highlighting its greater accuracy and efficiency. It could also be observed that when setting a higher $N_{c}$, the performance will not improve continuously, primarily because excessive corrections may confuse the agent and obscure correct answers.
MMLong TabBen MVision Average
No Reward 38.3 54.2 36.7 71.4
Agent-Specific Reward 42.8 56.3 40.5 72.7
Global Reward 34.1 51.6 35.1 70.2
No Scaling 34.9 50.8 34.5 71.1
Parallel Scaling 39.1 55.2 37.5 72.0
Sequential Scaling 41.3 54.7 40.0 72.4
Hybrid Scaling 39.8 55.5 39.4 73.0
Internal Scaling 41.8 55.9 41.6 72.3
Agent-Wise Hybrid 43.7 57.2 41.8 74.8
Scaling & Mixed Reward
Table 4: Results of different reward modeling and test-time scaling strategies.
(a) Analysis of Corrections
(b) Impact of $N_{p}$ and $N_{e}$
Figure 4: (a) Line graph shows the impact of various maximum numbers of corrections, with dashed and dotted lines denoting average values across all and three selected benchmarks, respectively. The bar charts show the average judgment numbers when the maximum is set to 3. (b) The line graphs represent the impacts of $N_{p}$ and $N_{e}$.
Analysis of Reward modeling and Test-Time Scaling
We perform experiments on our proposed mixed reward modeling and agent-wise hybrid test-time scaling. For the former, we separately use the agent-wise reward and global reward for comparison. For the latter, we benchmark our method with the other four effective test-time scaling strategies. Here, we select budget forcing (Muennighoff et al. 2025) as the internal scaling. Although the improvement from global reward is relatively limited, it avoids the selfishness of agents that only use agent-specific reward, yielding a further 1.3% increase in average scores. Our strategy, which designs tailored scaling forms for each agent according to its functions and specificity, outperforms all four existing strategies and improves the average scores by at least 1.8%. As demonstrated in Figure 4(b), higher values of $N_{p}$ and $N_{e}$ increase the likelihood of finding the correct answers, which leads to more attempts at plans and execution steps.
Conclusion
Our proposed MACT is an innovative multi-agent collaboration framework with a test-time scaling strategy for visual document understanding and question answering. Comprising four collaborative agents, this framework achieves effective functional role division through synergistic collaboration. Notably, the judgment agent strategy outperforms existing self-correction mechanisms while requiring fewer corrections. Additionally, mixed reward modeling and agent-wise hybrid scaling are particularly well-suited for multi-agent systems, as they strike a balance between agent-specific capabilities and collaborative efficiency. Extensive comparative experiments and supplementary analysis validate the superiority and effectiveness of our MACT. Furthermore, this work constitutes a meaningful exploration of multi-agent frameworks and test-time scaling strategies, underscoring promising directions for future research.
Appendix
Appendix A Related Works
General VLMs
Recent advancements in general VLMs have largely bridged the gap between visual and linguistic information, demonstrating impressive abilities in visual understanding, reasoning, and question answering. For example, latest GPT-4o (Hurst et al. 2024), GPT-4V (OpenAI 2024), Gemini-2.0-Pro (DeepMind 2025), DeepSeek-VL2 (Wu et al. 2024), Qwen2.5-VL (Bai et al. 2025), Claude-3.7-Sonnet (Anthropic 2024), InternVL-3 (Zhu et al. 2025), Llama-3.2-Vision (Grattafiori et al. 2024), Ovis2 (Lu et al. 2024b) and MiMo-VL (Xiaomi 2025) contribute to push the performance boundaries in many aspects. However, there is still room for improvement for these general VLMs in terms of specific downstream tasks in various scenarios. Furthermore, their derivative models, such as LLaVA-OneVision (Li et al. 2025b), Monkey (Li et al. 2024), Emu3 (Wang et al. 2024c), InternLM-XComposer2-4KHD (Dong et al. 2024a) and mPLUG-Owl3 (Ye et al. 2025), contribute to push the performance boundaries in many aspects.
Specialized VLMs
To better promote visual understanding and question answering, especially in document-based tasks, such as texts, webpages, charts, and tables, researchers have designed numerous task-specialized VLMs. For instance, TGDoc (Wang et al. 2023) and MMCA (Liu et al. 2024a) are explorations of document-based or chart-based understanding via large-scale collected instruction tuning data, and DocPedia (Feng et al. 2024) deals with the visual inputs at the frequency domain. To improve the comprehension of screenshots and webpages, UReader (Ye et al. 2023) designs a cropping strategy, while CogAgent (Hong et al. 2024) and TextMonkey (Liu et al. 2024b) use two resolution branches or a shifted window to accommodate document-oriented tasks. The series of works by mPLUG-DocOwl (Hu et al. 2024a, b) design a high-resolution compressor for multi-page document understanding with reduced visual token costs. Besides, ColPali (Faysse et al. 2025), M3DocRAG (Cho et al. 2024), and MDocAgent (Han et al. 2025) introduce efficient and novel retrieval mechanisms for VLVMs to comprehend long documents.
Multi-Agent Models
Agentic large models have gained tremendous popularity; however, single-agent systems are increasingly unable to meet the demands of complex and novel tasks. Thus, communicative agents’ mechanisms are implemented to improve cooperation and reaction between agents by two primary forms, i.e., debate (Du et al. 2024; Zhang et al. 2024a; Yue et al. 2025; Pan et al. 2024) and competition (Cheng et al. 2024; Chen et al. 2024a). Although the domain of LLM-based multi-agent models has expanded to include code generation (Islam, Ali, and Parvez 2024; Tao et al. 2024), financial decision-making (Yu et al. 2024b), and other areas, they still suffer from a lack of visual understanding. Thus, more recently, some recent works attempt to explore how multi-agent interaction benefits VLMs. For instance, Insight-V (Dong et al. 2024b) divides the functions on reasoning and summary agents, while Mobile-Agent-v2 (Wang et al. 2024a) and  (Li et al. 2025c) introduce specialized agents for planning. To accommodate different tasks, METAL (Li et al. 2025a), WebPilot (Zhang et al. 2025b), ViDoRAG (Wang et al. 2025b), MatPlotAgent (Yang et al. 2024), and VipAct (Zhang et al. 2024d) further break down the labor of a single agent into more manageable sub-instances and specialise the function of each role-tailored agent. Additionally, MuMA-ToM (Shi et al. 2025) and MDocAgent (Han et al. 2025) design separate image and text agents and then fuse them, ignoring the correlations between the two modalities. However, a practical multi-agent collaboration framework for visual document understanding and question answering is still deficient.
Test-Time Scaling
As a promising and burgeoning research focus, test-time scaling can further elicit various capabilities of models with relatively lower computational cost (Zhang et al. 2025a). The scaling could be categorized into four types: parallel scaling (Renze 2024; Brown et al. 2024; Wang et al. 2025a; Liu et al. 2025), generating multiple candidates and aggregating them; sequential scaling (Gou et al. 2024; Chen et al. 2024b; Yu et al. 2024a), iteratively updating intermediate states and triggering self-correction; Hybrid scaling (Besta et al. 2024; Guan et al. 2025; Snell et al. 2025) is another realization, which exploits their complementary benefits; internal scaling, which autonomously allocates computational tokens, e.g. budget forcing strategy (Muennighoff et al. 2025; Son et al. 2025; Li et al. 2025a). To further enhance the ability of models to scale at test time, supervised fine-tuning (SFT) and reinforcement learning (RL) are also widely employed to improve scaling performance. For example, an intuitive and efficient strategy is to set a reward model, such as process reward modeling (Snell et al. 2025; Zhao et al. 2025) and outcome reward modeling (Liu et al. 2025; Zuo et al. 2025; Setlur et al. 2025), offering quantitative reward values for test-time scaling.
Figure 5: Illustrations of representative samples from our selected 15 benchmarks on both document and non-docuement tpyes.
Appendix B Experimental Settings
Training Pipeline
In this work, we design a two-stage training pipeline to harness the full potential of our framework, specifically the SFT stage and RL stage. Moreover, we select three groups of base models for different variants, which are shown in Table 5. To enhance the general abilities, during the first stage, we employ AdamW optimizer with $\beta=\left(0.9,0.95\right)$, and the initial learning rate and weight decay are set to 2e52e{-5} and $0.1$ respectively. We train our model for one epoch with a batch size of 128. In the second RL stage, we adopt the GRPO (Shao et al. 2024) strategy for further optimisation. Following its original configurations, we set the learning rate, KL coefficient, and number of iterations as e6e{-6}, $0.04$, and $1$, respectively. The reward signals are generated by VisualPRM (Wang et al. 2025c) and Skywork-VL-Reward (Wang et al. 2025d); the former provides step-by-step signals for planning and execution agents, while the latter generates rewards for judgment and answer agents. During the process, we set the temperature parameter to $0.75$ to ensure diversity and creativity in the generated solutions. It should be noted that reward models and our framework are loaded on different servers, thereby avoiding the repeated loading and offloading of model weights. All the training and inference are conducted on 8 NVIDIA A100 GPUs.
Type Qwen2.5-VL Series MiMo-VL Series InternVL3 Series
(Bai et al. 2025; Yang et al. 2025) (Xiaomi 2025; Xia et al. 2025) (Zhu et al. 2025)
Planning Agent VLM Qwen2.5-VL-7B-Instruct MiMo-VL-7B-SFT InternVL3-9B-Instruct
Execution Agent VLM Qwen2.5-VL-7B-Instruct MiMo-VL-7B-SFT InternVL3-9B-Instruct
Judgment Agent LLM* Qwen2.5-7B-Instruct MiMo-7B-SFT InternVL3-8B-Instruct
Answer Agent LLM* Qwen2.5-3B-Instruct MiMo-7B-SFT InternVL3-3B-Instruct
Size - 7B+7B+7B+3B 7B+7B+7B+7B 9B+9B+8B+2B
=24B =28B =28B
Table 5: The base models of each variant of MACT. * indicates that we use VLMs, which remain powerful textual ability, because in InternVL3 series the model family does not provide LLMs.
Datasets
Our selected datasets cover four document types: text-based, webpage-based, table-based, and chart-based benchmarks, and also two non-document types: general and mathematical benchmarks. For each type, two to four datasets are included for both training and evaluation benchmarks. To better illustrate the datasets, statistical information provided in Table 6 and a brief introduction of each dataset are listed as follows:
Dataset Type Num I Num Q Avg Q Avg A Avg I Per Q
DocVQA (Mathew, Karatzas, and Jawahar 2021) Text 12K 50K 8.3 2.1 1.0
DUDE (Van Landeghem et al. 2023) Text 5K 41K 8.7 3.4 5.7
SlideVQA (Tanaka et al. 2023) Text 52K 15K 9.0 2.4 20.0
MMLongBench-Doc (Ma et al. 2025) Text 6K 1K 16.4 2.8 47.5
VisualMRC (Tanaka, Nishida, and Yoshida 2021) Webpage 10K 30K 10.6 9.5 1.0
InfographicVQA (Mathew et al. 2022) Webpage 5K 30K 11.5 1.6 1.2
ChartQA (Masry et al. 2022) Chart 22K 34K 7.2 1.2 1.0
CharXiv (Wang et al. 2024d) Chart 2K 13K 20.4 2.9 1.0
TableVQA-Bench (Kim, Yim, and Song 2024) Table 1K 2K 10.5 4.7 1.0
TableBench (Wu et al. 2025) Table 21K 4K 20.3 8.5 1.0
ScienceQA (Saikh et al. 2022) General 10K 21K 12.1 4.4 0.5
RealWorldQA (xAI 2024) General 1K 1K 11.1 1.2 1.0
MathVista (Lu et al. 2024a) Mathematical 5K 6K 15.6 1.2 1.0
Math-Vision (Wang et al. 2024b) Mathematical 3K 3K 42.3 2.2 1.0
MathVerse (Zhang et al. 2024c) Mathematical 3K 16K 35.7 1.4 1.0
Table 6: Statistical information of our selected datasets. Here, Num and Avg indicate the total number and the average length, while I, Q, and A represent images, questions, and answers, respectively.
  • DocVQA (Mathew, Karatzas, and Jawahar 2021). It is a classic document-based dataset, with more than 12K images from various documents and 50K related questions. Each question corresponds to a single-page document, which requires relatively simple visual understanding. The average lengths of questions and answers are 8.3 and 2.1 words.
  • DUDE (Van Landeghem et al. 2023). It is a multi-page document question answering dataset, comprising 41,541 questions and over 5,000 documents. The average number of pages in the document of this dataset is 5.72, while the average lengths of the questions and answers are 8.7 and 3.4, respectively.
  • SlideVQA (Tanaka et al. 2023). It is a slide question answering dataset, which incorporates 52K images from slides and about 15K questions. The average lengths of the questions and answers are 9.0 and 2.4 words, and they include a certain proportion of multi-hop and numerical reasoning.
  • MMLongBench-Doc (Ma et al. 2025). It introduces a multi-page dataset for document understanding, consisting of about 6K document images and 1K questions. Depending on multi-hop evidence, the questions and answers have 16.4 and 2.8 words on average, and almost half of the questions are cross-page, while one-fifth of the questions are unanswerable.
  • VisualMRC (Tanaka, Nishida, and Yoshida 2021). It collects 10K images sourced from 35 domains of webpages and more than 30K annotated questions with an average length of 10.6 words. The answers, with an average length of 9.5 words, in this dataset include both word spans in the context and the layouts of the webpages, which are more complex and diverse, with more abstract questions.
  • InfographicVQA (Mathew et al. 2022). It consists of more than 5K images collected from webpages and 30K questions, constructed with textual, graphical, and visual elements. Most questions in this dataset are extractive and numerical, testing basic reasoning and arithmetic skills, with 11.5 and 1.6 words in questions and answers on average.
  • ChartQA (Masry et al. 2022). It introduces a dataset with 22K images and 34K questions, involving bar, line, pie charts, and composite plots. The questions in this dataset require visual, logical, and arithmetic ability, with an average length of questions and answers of 7.2 and 1.2 words.
  • CharXiv (Wang et al. 2024d). It presents a dataset for chart understanding and answering, with 2K chart images and 13K curated questions. It includes both descriptive and reasoning sub-datasets, and we merge them. The questions and answers have 22.6 and 2.8 words evenly.
  • TableBench (Wu et al. 2025). It is a dataset with tabular and textual hybrid contents, constructed with over 4K tables and 21K questions, with numerical reasoning, such as computing, comparison, and sorting. Its questions and answers have lengths of 20.3 and 8.5 words on average.
  • TableVQA-Bench (Kim, Yim, and Song 2024). It establishes a dataset for table-based question answering, extended with four existing sources with 1K tables and 2K questions. Consisting of diverse types of questions, the average lengths of questions and answers are 10.5 and 4.7 words.
  • ScienceQA (Saikh et al. 2022). It is a dataset that includes about 21K multiple-choice questions based on more than 10K images from diverse scientific fields, and the length of the questions and answers is 12.1 words on average.
  • RealWorldQA (xAI 2024). It consists of about 1K questions and 1K images. The average length of the questions in this dataset is 11.1 words, while that of the answers is only 1.2 words.
  • MathVista (Lu et al. 2024a). It proposed a dataset integrating diverse mathematical and visual tasks with 5K images and 6K questions, demanding fine-grained visual understanding and compositional reasoning. On average, the length of the questions and answers is 15.6 and 1.2 words, respectively.
  • Math-Vision (Wang et al. 2024b). It is a dataset integrating diverse mathematical tasks with about 3K images and 3K questions from real exams, spanning 16 distinct mathematical disciplines and graded across 5 levels of difficulty. It has the longest average length of questions (42.3 words) in our selected datasets, and 2.2 words in answers on average.
  • MathVerse (Laurençon et al. 2024). It is a mathematical dataset with about 3K images and 16K questions. On average, the lengths of the questions and answers in this dataset are 35.7 and 1.4 words, respectively.
For some datasets, we use pre-treatments to handle the visual inputs. For the document in some datasets, which is provided in PDF format, we transform these files into visual images as input. Besides, we design a standardised visualisation process for part of the tables in the TableBench (Wu et al. 2025) dataset, which only contains HTML format data without authentic visual images. Additionally, we utilize the PDF2Image library to convert PDF files into images for input.
Evaluations
We onboard the unavailable models in LMMs-Eval (Zhang et al. 2024b) to evaluate their performances with the consistent configurations. We directly use the natively supported benchmarks in the LMMs-Eval framework, namely DocVQA (Mathew, Karatzas, and Jawahar 2021), InfographicVQA (Mathew et al. 2022), ChartQA (Masry et al. 2022), ScienceQA (Saikh et al. 2022), RealWorldQA (xAI 2024), MathVista (Lu et al. 2024a), and MathVerse (Zhang et al. 2024c). In terms of unavailable benchmarks in LMMs-Eval, namely DUDE (Van Landeghem et al. 2023), SlideVQA (Tanaka et al. 2023), MMLongBench-Doc (Ma et al. 2025), VisualMRC (Tanaka, Nishida, and Yoshida 2021), CharXiv (Wang et al. 2024d), TableVQA-Bench (Kim, Yim, and Song 2024), TableBench (Wu et al. 2025), and Math-Vision (Wang et al. 2024b), we onboard the benchmarks and register them by YAML configurations.
To evaluate the performance of models on these benchmarks fairly and equitably, we employ a range of evaluation metrics. For DocVQA (Mathew, Karatzas, and Jawahar 2021), DUDE (Van Landeghem et al. 2023), InfographicVQA (Mathew et al. 2022) benchmarks, we follow the original evaluation metrics, i.e., ANLS (Average Normalised Levenshtein Similarity) (Biten et al. 2019), to measure the difference between the predicted answers and ground truths. Moreover, the F1 score is utilized for SlideVQA (Tanaka et al. 2023). For other selected benchmarks, we use powerful GPT-4o (Hurst et al. 2024) as a judge model to evaluate the answers by LMMs-Eval, which generates scores from 1 to 10, and then the scores will be scaled to an accuracy between 0 and 100 in the dimension of benchmarks.
Figure 6: Prompts for the four agents.
Prompt Templates
As shown in Figure 6, there are prompts for the four agents, including the system prompts and the prompts for each process. Apart from the system prompts, $\mathcal{A}_{plan}$ and $\mathcal{A}_{exe}$ have two extra prompts to accomplish their functions, while $\mathcal{A}_{judg}$ and $\mathcal{A}_{ans}$ have only one. In the separating prompt of $\mathcal{A}_{exe}$, each execution unit is comprised of: (1) a specific definition; (2) expected target and output; (3) existing inputs or results from the previous step.
Appendix C Results and Discussions
Additional Quantitative Results
For the line graphs and bar charts shown in Figure 4(a) and the line graphs in Figure 4(b), we list the quantitative results in Table 7 and Table 8, respectively, for clarity. Compared with the baseline, our judgment agent strategy for correction improves the scores by 6.2% on average when the maximum number of corrections is set to 3. Furthermore, it still has 5.3% and 2.6% increases when compared to the other two correction strategies. Besides, our strategy achieves the best performance with fewer average number of corrections, indicating high efficiency and accuracy. In terms of the impact of the number of generated plans $N_{p}$ and the number of candidate executions $N_{e}$, it is evident that higher values of the two parameters cause superior performance.
Max Num of Corrections Avg Num of
3 5 7 10 Corrections
Baseline 40.2 -
Internal Correction 40.8 42.0 41.5 38.2 2.8
Agent for Judgment and Correction 45.1 46.7 44.6 41.2 2.3
Independent Judgment Agent (Ours) 47.6 47.5 45.3 42.9 1.9
Baseline 68.8 -
Internal Correction 69.5 70.3 69.6 64.2 2.3
Agent for Judgment and Correction 72.2 72.6 71.9 68.9 1.6
Independent Judgment Agent (Ours) 74.8 74.7 73.7 71.2 1.3
Table 7: Comparisons among the three correction strategies of various maximum numbers of corrections and the average number of corrections when the limitation is set to 3. Line with denotes the results on the three selected benchmarks, while another line is the average results on all benchmarks.
$N_{p}$ $N_{e}$ 1 2 4 8 16
1 70.2 70.8 72.0 72.5 72.8
2 71.1 - - - 73.4
4 71.5 - - - 73.9
8 72.0 - - - 74.5
16 72.2 72.9 73.4 74.1 74.8
Table 8: Results of different values of $N_{p}$ and $N_{e}$ on all benchmarks.

### v2

@@
-Visual Document Understanding and Question Answering: A Multi-Agent Collaboration Framework with Test-Time Scaling
+Visual Document Understanding and Reasoning: A Multi-Agent Collaboration Framework with Agent-Wise Adaptive Test-Time Scaling
 Abstract
-Existing vision-language models (VLMs), whether generalists or specialists, remain constrained by their parameter scale, lack robust self-correction capabilities, and underperform in tasks involving long visual contexts and complex reasoning, resulting in suboptimal performance on document-based tasks.
-To address this, we propose MACT, a Multi-Agent Collaboration framework with Test-Time scaling, tailored for visual document understanding and visual question answering (VQA).
-It comprises four distinct small-scale agents, i.e., planning, execution, judgment, and answer agents, with clearly defined roles and effective collaboration.
-Notably, the judgment agent exclusively verifies correctness and redirects to prior agents for revisions, outperforming conventional correction strategies.
-To further expand the capability boundaries of the framework, we propose mixed reward modeling that balances agent-specific abilities and global collaboration, as well as agent-wise hybrid test-time scaling, which customizes different scaling strategies for each agent based on their functions.
-Evaluated on benchmarks spanning both document-based and non-document-based settings, our MACT shows superior performance with a smaller parameter scale without sacrificing the ability of general and mathematical tasks.
-Especially, it stands out in benchmarks involving long visual contexts and complicated reasoning.
-The three variants of MACT consistently hold the top three positions in average scores, leading in 13 of the 15 benchmarks.
-Code will be available at: https://github.com/YU-deep/MACT.git.
-Introduction
-For humans, acquiring and understanding visual information from documents—such as texts, webpages, charts, and tables—is a common and indispensable part of daily life.
-As a derivative task, document-based VQA provides an assessment of the understanding of both textual and visual document information as well as their relations, whose performance is quantifiable.
-Therefore, numerous studies aim to propose benchmarks for more complex document-based tasks of diverse types to evaluate model performance fairly and comprehensively.
-Specifically, the types have expanded from single-page documents, simple charts, and tables to real-world multi-page documents, complicated webpages, bar charts, pie charts, line charts, composite tables, matrix tables, and mathematical reasoning.
-Additionally, the form of the questions has also transformed from straightforward Yes/No, single-choice or direct information extraction questions to diversified question types, including but not limited to visual details detection, cross-page or multi-hop information extraction and summarization, comparison, multi-step calculation, intersection finding and unanswerable questions, which are answered based on the layout, structure, visual and textual contents.
-Thus, the latest flagship VLMs of various series are dedicated to expanding these boundaries of understanding documents.
-(a) Qwen2.5-VL Series Based
-(b) MiMo-VL Series Based
-(c) InternVL-3 Series Based
-Figure 1: Comparisons among the three variants of our MACT, their base models, and larger models in the same series.
-Recently, the latest released general VLMs have showcased exceptional understanding and generalization capabilities.
-For instance, closed-source models, such as GPT-4o (Hurst et al.
-2024), Gemini-2.0-Pro (DeepMind 2025) and Claude-3.7-Sonnet (Anthropic 2024) have surpassed human-level scores in certain tasks, while open-source models, e.g., Llama-3.2-Vision (Grattafiori et al.
-2024), Qwen2.5-VL (Bai et al.
-2025), MiMo-VL (Xia et al.
-2025), InternVL-3 (Zhu et al.
-2025) and Ovis2 (Lu et al.
-2024b) have also achieved great performance.
-However, their performances in document-based tasks remain suboptimal due to the biases or deficiencies in the training data.
-To further enhance the ability, some researchers have endeavored to develop specialized VLMs.
-For example, UReader (Ye et al.
-2023), TextMonkey (Liu et al.
-2024b) and mPLUG-DocOwl2 (Hu et al.
-2024b) focus on document understanding.
-However, their overall performance lags far behind that of the latest general VLMs with larger parameters.
-To summarize, currently existing VLMs, including both generalists and specialists, exhibit three key limitations in understanding visual documents and answering questions: (1) Constrained by Parameter Scale: There are substantial performance gaps between larger-parameter VLMs and smaller counterparts in the document domain, and the potential of smaller-parameter VLMs remains underactivated.
-(2) Deficiency in Self-Correction Capacity: As self-correction is indispensable when humans tackle tasks involving challenging and complex documents, existing methods either fail to adequately address this capability or adopt suboptimal designs.
-(3) Underperformance on Long Visual Contexts and Reasoning: Results on some document-based benchmarks, which feature long visual contexts or require dense reasoning, are unsatisfactory, with notably low accuracy.
-Given these issues, we propose a multi-agent collaboration framework with a dedicated test-time scaling strategy for visual document understanding and question answering.
-Within this framework, we employ four relatively lightweight collaborative agents: planning, execution, judgment, and answer agents.
-Through their collaboration and mixed reward modeling, each agent directs attention to both agent-tailored and global objectives, thereby unleashing the potential of the agents with small-scale parameters.
-For more efficient self-correction capacity, we design a mechanism that employs an additional judgment agent.
-It only checks the correctness of the generated execution plans and execution processes, with corrections implemented by the preceding agents.
-Additionally, we introduce an agent-wise hybrid test-time scaling approach that considers the distinct functions and features of each agent, markedly improving abilities in handling long visual contexts and complex reasoning.
-We train three variants of MACT based on different groups of base models using a two-stage pipeline, and evaluate them on 15 benchmarks, comprising 10 document-based and 5 non-document-based ones, to comprehensively assess document understanding as well as general and reasoning performance.
-Compared to models of comparable scale and larger-scale ones, the results indicate that the three variants of MACT show an average increase of at least 3.2% to 5.6%, securing the top three positions and achieving the best performance in 13 out of 15 benchmarks.
-Furthermore, our models have distinct advantages in long visual contexts and reasoning tasks, as further validated by results on selected non-document benchmarks.
-Methodology
+The dominant paradigm of monolithic scaling in Vision-Language Models (VLMs) is failing for understanding and reasoning in documents, yielding diminishing returns as it struggles with the inherent need of this domain for document-based procedural reasoning, cognitive complexity, and factual accuracy.
+To this end, we introduce MACT, a Multi-Agent Collaboration framework with agent-wise adaptive Test-time scaling that pioneers a paradigm shift to procedural scaling, adapting dynamically to the functional entities of visual documents understanding and reasoning.
+MACT decomposes the visual document processing flow into four specialized agents, i.e., planning, execution, judgment, and answer, to resolve cognitive overload and introduce a critical self-correction loop for factual grounding.
+This collaborative architecture is amplified by an agent-wise adaptive test-time scaling strategy that intelligently allocates computational resources based on the complexity and redundancy of each functionality.
+Evaluated on multiple visual document understanding benchmarks, MACT achieves superior performance with a smaller parameter scale, adapting effectively to various document scenarios without compromising its general or mathematical reasoning capabilities.
+The three variants of MACT consistently attain top-three average performance rankings, with average performance enhancements of 9.9–11.5% over the base models.
+The source code will be released publicly.
+1 Introduction
+For humans, acquiring and understanding visual information from documents is an indispensable part of daily life.
+The corresponding tasks of visual documents, e.g., Visual Question Answering (VQA), serves as a quantifiable benchmark for a model’s ability to comprehend complex layouts, dense text, and structured data.
+To this end, the field has seen a rapid evolution of benchmarks, moving from simple documents to multi-page reports, complex charts, and webpages.
+The questions have similarly evolved from simple lookups to requiring semantic understanding and multi-hop reasoning, pushing the capability boundaries of VLMs.
+(a) Comparisons on Qwen2.5-VL Series [qwen2025qwen25technicalreport]
+(b) Comparisons on MiMo-VL Series [xia2025mimo]
+(c) Comparisons on InternVL-3 Series [zhu2025internvl3]
+Figure 1: Comparisons among the three variants of MACT, the base models of these variants, and larger-scale models within the same model family, indicating the superiority of our framework over monolithic scaling paradigms.
+The dominant paradigm for advancing VLM capabilities, both for documents and other general vision data, has been monolithic scaling, i.e., increasing parameter size, leveraging more high-quality training data.
+The approach has led to remarkable success, with phenomenal models like GPT-4o [hurst2024gpt], Gemini-2.0-Pro [gemini2024deepmind] and Claude-3.7-Sonnet [anthropic2024claude], demonstrating powerful general-purpose abilities.
+However, for document-based VQA, this strategy of uniform scaling is showing diminishing returns.
+As shown in Fig.
+1, only limited performance gain is attained along parameter scaling for open-sourced models [qwen2025qwen25technicalreport, zhu2025internvl3] with exponential growth in computational overhead, suggesting that simply increasing parameter count is a brute-force solution that fails to address the core challenges of the domain.
+The core problem is that documents are not monolithic entities that are suitable for uniform scaling.
+Unlike natural images, they possess a unique set of characteristics that make naive scaling less effective: (1) Suboptimal procedural reasoning.
+Document analysis is not a single action of perception but rather an inherently procedural, multi-step workflow [kutsyuruba2023document]: decomposing the question, forming a strategy, locating relevant information, and synthesizing a final answer.
+Monolithic models, with their single forward-pass architecture, attempt to solve this entire procedure at once, leading to less robust reasoning paths.
+(2) Cognitive overload.
+The procedural steps of document analysis demand a diverse set of specialized skills, such as global layout parsing, fine-grained text extraction, logical inference, and numerical calculation.
+A monolithic model is forced to encode all these disparate functions into a single set of weights, leading to cognitive overload and task interference.
+This leads to a compromise where the model achieves broad generalization at the expense of deep expertise in any single task.
+(3) Vulnerability to factual errors.
+The semantic meaning in documents is extremely sensitive to minor procedural mistakes [fitzgerald2012documents], even paragraph truncation errors or misaligning a table row can invalidate the entire answer.
+Monolithic, feed-forward models lack an internal verification or self-correction loop.
+An initial extraction error, however small, inevitably cascades through the reasoning process without being challenged, identified, and corrected.
+These challenges reveal the need for a new paradigm: a shift from monolithic scaling to procedural scaling, where the problem is decomposed and specialize test-time scalings are applied to each entity.
+To achieve this, we propose MACT, a Multi-Agent Collaboration framework with agent-wise adaptive Test-time scaling, enabling a more effective scaling framework.
+Specifically, it deconstructs the monolithic model into four specialized agents that explicitly mirror the required cognitive functions.
+Within this framework, we employ four relatively collaborative agents: planning, execution, judgment, and answer agents.
+This design inherently enforces procedural reasoning, dedicating a specialist to each cognitive step from planning to synthesis.
+This decomposition is what resolves the cognitive overload problem, allowing each agent to achieve mastery over its specific function.
+The key to unlocking this specialized performance is our agent-wise adaptive test-time scaling approach.
+Rather than uniform scaling, it is intelligent, on-demand allocation of computational resources, allowing the execution agent to apply maximum scrutiny to a single data point while the planning agent operates efficiently on the whole document.
+Finally, the framework directly confronts the high cost of errors.
+The judgment agent acts as a built-in verification, providing the self-correction capabilities that monolithic models lack.
+We train three variants of MACT based on different groups of base models using a two-stage pipeline, and evaluate them on 15 benchmarks, comprising 10 document-based and 5 non-document-based ones.
+Our MACT showcases superior document understanding and reasoning without compromising general capabilities.
+Compared to models of comparable scale and larger-scale ones, the three variants of MACT demonstrate an average increase of 3.2% to 5.6%, securing the top three positions and achieving the best performance in 13 out of 15 benchmarks.
+The results demonstrate that monolithic scaling is not the optimal approach for visual document-based tasks.
+In contrast, the adaptive test-time scaling in multi-agent framework emerges as a more advantageous alternative, as it achieves superior scaling performance.
+2 Methodology
 This section provides a detailed elaboration of our proposed MACT, divided into three subsections.
-We first formally define the task, then introduce the four agent components: the planning agent, the execution agent, the judgment agent, and the answer agent, as well as the collaboration mechanisms among them.
-Finally, we detail the mixed reward modeling and agent-wise hybrid test-time scaling tailored to our multi-agent collaboration framework.
-Task Definition
-For the question answering task defined in this work, given a question $\mathcal{Q}$ and the its corresponding visual inputs $\mathcal{V}=\{v_{1},v_{2},\dots,v_{n}\}$, the target is to generate a correct answer output $\mathcal{O}$.
-This answer should leverage the information within $\mathcal{V}$ and answer the proposed question comprehensively and accurately.
+We first introduce the multi-agent framework, which decomposes the document-tailored workflow into four functionalities: planning, execution, judgment, and answer, and defines the collaboration mechanisms among them.
+Then, we detail the agent-wise adaptive test-time scaling and mixed reward modeling tailored to the multi-agent collaboration framework.
+More details are presented in Appendix 7.
 Figure 2: The overview of MACT.
-The upper part demonstrates our proposed multi-agent framework, which has four collaborative agents for tailored tasks.
+The upper part demonstrates our procedural framework, with four tailored and collaborative agents to conduct the process of document analysis.
 When the judgment agent detects mistakes, it redirects to previous agents for corrections.
-While the lower part illustrates the mixed reward modeling and agent-wise hybrid scaling for the multi-agent framework.
-Multi-Agent Collaboration Framework
-As shown in the upper part of Figure 2, MACT consists of four collaborative agents: sequentially planning agent ($\mathcal{A}_{plan}$), execution agent ($\mathcal{A}_{exe}$), judgment agent ($\mathcal{A}_{judg}$), and answer agent ($\mathcal{A}_{ans}$).
-Each agent is role-specialized, which accomplishes its functionality and then passes the generated contents to the subsequent agent or outputs the final answers.
+The lower part illustrates the agent-wise adaptive test-time scaling and mixed reward modeling for the multi-agent framework.
+2.1 Multi-Agent Collaboration Framework
+Document understanding and reasoning constitutes not merely a unilateral perceptual activity, but an expected multistage process [kutsyuruba2023document]: decomposing the question, formulating execution strategies, locating information pertinent to the document, conducting validation with necessary corrections, and synthesizing a final response.
+As shown in the upper part of Fig.
+2, MACT consists of four collaborative agents: sequentially planning agent ($\mathcal{A}_{plan}$), execution agent ($\mathcal{A}_{exe}$), judgment agent ($\mathcal{A}_{judg}$), and answer agent ($\mathcal{A}_{ans}$).
+Each agent is role-specialized, which accomplishes its functionality and then passes to the subsequent agent or outputs the final answers.
 [Uncaptioned image] Planning Agent.
-$\mathcal{A}_{plan}$ mainly focuses on the analysis and decomposition of the original question, making high-level executive plans.
-Largely following the analogical prompting principles (Yasunaga et al.
-2024), it first generates $\mathit{N_{p}}$ relevant sample problems along with their corresponding plans simultaneously.
-By referring to each sample problem and its corresponding plan, it produces distinct execution plans for the current problem respectively, providing diverse pathways to accomplishment.
-It should be noted that this agent only yields high-level execution plans without implementation details.
-It ensures that $\mathcal{A}_{plan}$ can better formulate the plan from a holistic perspective, and avoid interfering with the execution of $\mathcal{A}_{exe}$.
-Moreover, it avoids premature binding to specific tools, enabling $\mathcal{A}_{exe}$ to select optimal resources for diverse tasks dynamically.
+Given a question $Q$ and corresponding visual document inputs $\mathcal{D}$, $\mathcal{A}_{plan}$ mainly focuses on the analysis and decomposition of the original question, making high-level executive plans on the document-based tasks.
+Before generating the plans for the original task, we generate $\mathit{N_{p}}$ similar and relevant instances, and their plans simultaneously, which is adapted from analogical prompting principles [michihiro2024large].
+For each generated example and corresponding plan, $\mathcal{A}_{plan}$ generates a distinct execution plan respectively, which could be formulated as:
+$\mathcal{P}_{rel}$ $=\mathcal{A}_{\text{plan}}\left(\mathcal{Q},\mathcal{D}\right),$ (1)
+$\mathcal{P}$ $=\mathcal{A}_{\text{plan}}\left(\mathcal{Q},\mathcal{P}_{rel},\mathcal{D},\mathcal{M}\right),$ (2)
+where generated relevant plans $\mathcal{P}_{rel}=\left\{r_{1},r_{2},\dots,r_{\mathit{N_{p}}}\right\}$, and execution plans $\mathcal{P}=\left\{p_{1},p_{2},\dots,p_{\mathit{N_{p}}}\right\}$, which provide different pathways to accomplishment.
+Besides, $\mathcal{M}$ is the mistake from $\mathcal{A}_{judg}$ for correction, and it is initialized to empty.
+For each relevant plan and execution plan $p$, it composed of steps $\left\{s_{1},s_{2},\dots,s_{n}\right\}$.
+It should be noted that we only yield high-level execution plans with limited details in this agent.
+To be more precise, the high-level plan describes the expected targets and requirements of each step, but does not output specific execution processes directly.
+It ensures that $\mathcal{A}_{plan}$ can better project the overall plan instead of execution details and would not interfere with the execution of $\mathcal{A}_{exe}$.
+Besides, it avoids premature commitment to specific tools, enabling $\mathcal{A}_{exe}$ to dynamically select optimal resources for diverse tasks.
 [Uncaptioned image] Execution Agent.
-$\mathcal{A}_{exe}$ is designed to execute plans step by step and output the execution process using the selected tools from the tool library.
-Specifically, it breaks down the steps in the plan into execution units, with information for each unit populated into a template.
-$\mathcal{A}_{exe}$ then executes these units sequentially, deriving the intended outputs from each.
-Once all steps are completed, it concatenates the full output and the execution processes of all steps in sequence, passing them to subsequent agents.
+$\mathcal{A}_{exe}$ is designed to execute plans step by step and output the execution process using the selected tools from the tool library $\mathcal{T}=\{t_{1},t_{2},\dots,t_{n}\}$ (see in Appendix 7.2).
+Specifically, it breaks down the execution plan and each step will be regarded as an execution unit, with information for each unit populated into a template and comprised of: (1) a specific definition; (2) expected target and output; (3) existing inputs or results from the previous step.
+$\mathcal{A}_{exe}$ then executes the units sequentially, deriving the intended outputs from each.
+The execution of single step could be presented as:
+$e_{i}=\mathcal{A}_{\text{exe}}\!\left(\mathcal{Q},\mathcal{D},s_{i},\mathcal{T},\mathcal{M}\right).$ (3)
+Once all steps are completed, it concatenates the full output and the execution processes of all steps in sequence, passing $\mathcal{E}=\left\{e_{1},e_{2},\dots,e_{n}\right\}$ to subsequent agents.
 [Uncaptioned image] Judgment Agent.
-To enhance the quality of completing challenging tasks, humans often judge and correct mistakes that arise in planning or execution—such as reversed sequences, incorrect dependencies, redundant or missing steps, mismatches between plans and processes, and disconnects between steps.
-Similarly, some multi-agent systems, primarily focused on code generation, incorporate mechanisms that either (a) internally correct errors within the same agent (Yang et al.
-2024; Sun et al.
-2024) or (b) deploy an additional agent to handle both judgment and correction (Islam, Ali, and Parvez 2024; Li et al.
-2025a).
-However, the internal correction mechanism (a) struggles to identify most mistakes and may fall into cognitive blind spots, as generation and correction rely on the same model.
+Document-based visual tasks exhibit high sensitivity to process-related errors [fitzgerald2012documents], especially in monolithic models that omit verification and correction procedures.
+Trivial procedural discrepancies inevitably distort the overall reasoning trajectory, leading to a snowball effect that results in wholly erroneous outputs.
+Thus, some multi-agent systems incorporate mechanisms that either (a) internally correct errors within the same agent [yang2024matplotagent, sun2024pearl] or (b) deploy an additional agent to handle both judgment and correction [islam2024mapcoder, li2025metal].
+However, the (a) internal correction mechanism struggles to identify most mistakes and may fall into cognitive blind spots, as generation and correction rely on the same model.
 In contrast, the latter approach (b) requires the agent to possess strong capabilities for both judgment and correction, thereby necessitating models with larger parameters and more complex reward modeling designs.
 Furthermore, using a separate agent to regenerate mistaken components may result in incompatibilities or even conflicts with existing parts, given the autonomy of the two agents.
-In RL, the optimization objective for both (a) and (b), as illustrated in Figure 3, is to pass verification.
+In RL, the optimization objective for both (a) and (b), as illustrated in Fig.
+3, is to pass verification.
 This can lead to strategic production of vague statements or the omission of details, resulting in corrections that appear correct superficially but are actually misleading.
-To address the limitations of both approaches, we design a novel judgment strategy for multi-agent systems, which introduces an additional judgment agent to separate judgment from correction, thereby fostering specialization of labor.
+And utilizing another agent to regenerate the mistaken part might be incompatible or even causing conflicts with the existing parts because of the independence of two autonomous agents.
+To address the limitations of both approaches, we design a simple but effective judgment strategy for multi-agent systems, which introduces an additional judgment agent to separate judgment from correction, thereby fostering specialization of labor.
 More specifically, $\mathcal{A}_{judg}$ is primarily responsible for assessing the correctness of steps in previously generated execution plans and processes, without engaging in direct correction.
-If mistakes are detected in any step, $\mathcal{A}_{judg}$ identifies the specific problematic step, provides a brief description of the error, and routes the issue to the appropriate agent—either $\mathcal{A}_{plan}$ or $\mathcal{A}_{exe}$—for correction.
+The judgment process is summarized as:
+$\mathcal{J}=\mathcal{A}_{judg}\left(\mathcal{Q},\mathcal{P},\mathcal{E}\right),$ (4)
+where $\mathcal{J}=\left\{flag_{plan},flag_{exe},\mathcal{M}\right\}$, $flag_{plan}=flag_{exe}\in\left\{true,false\right\}$, and $\mathcal{M}$ is the mistake description.
+If mistakes are detected in any step, $\mathcal{A}_{judg}$ identifies the specific problematic step, provides a brief description of the mistake, and routes the issue to the appropriate agent, either $\mathcal{A}_{plan}$ or $\mathcal{A}_{exe}$, for correction.
 For each plan-process pair that is already correct or has been corrected, $\mathcal{A}_{judg}$ forwards them to $\mathcal{A}_{ans}$.
@@
 [Uncaptioned image] Answer Agent.
+The primary function of $\mathcal{A}_{ans}$ is to output the final answer through both the incorrect processes and the correct process.
 Counterintuitively, $\mathcal{A}_{ans}$ incorporates both the correct execution process and incorrect segments from prior processes.
 This facilitates direct focus on modifications within the corrected process, thereby preventing the omission of error-prone details.
+Additionally, the mistake-correction pair presents a natural and complete logical closed-loop structure, which benefits answer generation based on the execution process.
+Thus, the final output answer is:
+$\mathcal{O}=\mathcal{A}_{ans}\left(\mathcal{Q},\mathcal{E},\mathcal{M}\right).$ (5)
 Collaboration.
@@
 Algorithm 1 Multi-Agent Collaboration Procedure
-0: question $\mathcal{Q}$, visual inputs $\mathcal{V}=\{v_{1},v_{2},\dots,v_{n}\}$, planning agent $\mathcal{A}_{plan}$, execution agent $\mathcal{A}_{exe}$, the tool library of the former agent $\mathcal{T}=\{t_{1},t_{2},\dots,t_{n}\}$, judgment agent $\mathcal{A}_{judg}$, answer agent $\mathcal{A}_{ans}$, maximum number of corrections $N_{c}$
+0: question $\mathcal{Q}$, visual document inputs $\mathcal{D}=\{v_{1},v_{2},\dots,v_{n}\}$, planning agent $\mathcal{A}_{plan}$, execution agent $\mathcal{A}_{exe}$, the tool library of the former agent $\mathcal{T}=\{t_{1},t_{2},\dots,t_{n}\}$, judgment agent $\mathcal{A}_{judg}$, answer agent $\mathcal{A}_{ans}$, maximum number of corrections $N_{c}$
 0: answer output to the question $\mathcal{O}$
@@
 […]
-8:  $p_{2}\leftarrow\mathcal{PF}\left(\mathcal{Q},\mathcal{P}_{relevant},\mathcal{M}\right)$
+8:  $p_{2}\leftarrow\mathcal{PF}\left(\mathcal{Q},\mathcal{P}_{rel},\mathcal{M}\right)$
 […]
@@
 […]
-Mixed Reward modeling
-To guide RL in our multi-agent system and further enhance both the specialized and collaborative capabilities of each agent, we design a mixed reward strategy that integrates agent-specific and global outcome reward signals.
-Given that the four agents in the framework have distinct functions and reward signal preferences, we first incorporate agent-tailored rewards for each.
-In particular, for $\mathcal{A}_{plan}$ and $\mathcal{A}_{exe}$, whose outputs are step-by-step, we employ an effective multi-modal reward model to yield step-wise process reward signals.
-These step-wise rewards provide instant feedback for each step and hierarchical rewards for the two agents, boosting accuracy on challenging tasks.
-For $\mathcal{A}_{judg}$ and $\mathcal{A}_{ans}$, we directly use a reward model to generate one reward signal for each output of each agent.
-Furthermore, we apply a global reward, computed based on the final selected answer along the chosen path of the four agents.
-It strengthens rewards for correct paths, guiding the model to explore valid paths while mitigating adverse effects from erroneous ones.
-Besides, it mitigates the selfishness of agents to some extent, as individual agents tend to maximize their own rewards without regard for the global objective, and this global reward helps to alleviate such behavior.
-Agent-Wise Hybrid Test-Time Scaling
-Since that scaling test-time computation can activate the long-context understanding and reasoning capabilities of small-parameter VLMs, dynamically allocating additional computing resources and enabling multiple paths in the test time, we propose an agent-wise hybrid scaling strategy during inference.
+2.2 Agent-Wise Adaptive Test-Time Scaling
+For visual document understanding and reasoning, the key challenge is not visual documents themselves, but the procedural reasoning that links global question decomposition, plan formation, information extraction, and mistake judgment across multiple steps.
+Monolithic scaling enlarges model capacity uniformly, yet fails to accommodate this structured reasoning process.
+We therefore propose an agent-wise adaptive scaling strategy during inference, dynamically allocating additional computing resources to different functional entities and enabling on-demand allocation during the test time.
 Existing test-time scaling methods fall into four main categories: parallel scaling, sequential scaling, hybrid scaling, and internal scaling.
 However, these strategies are originally designed for single models or agents, overlooking the different division of labor among agents and achieving suboptimal performance when applied to multi-agent systems.
-Accordingly, we propose an agent-wise hybrid scaling strategy tailored to multi-agent architectures, which significantly enhances both agent-specific capabilities and collaborative performance.
-For the first three agents with various functions, we customize different test-time scaling strategies to their unique characteristics: (1) Planning Agent: In our design of $\mathcal{A}_{plan}$, it naturally provides relevant sample plans as references for formulating multiple execution plans for the original task.
-We therefore prompt the generation of $N_{p}$ relevant plans independently, yielding $N_{p}$ parallel paths per query.
-It establishes a foundation for scaling subsequent agents and increases the likelihood that at least one path will produce the correct answer.
-(2) Execution Agent: Given that the execution process is divided by step in $\mathcal{A}_{exe}$, we treat each step as an evaluation node.
+Accordingly, we propose an agent-wise adaptive scaling strategy tailored to multi-agent architectures, which significantly enhances both agent-specific capabilities and collaborative performance.
+For the first three agents with various functions, we customize different test-time scaling strategies to their unique characteristics: (1) [Uncaptioned image] Planning agent: In our design of $\mathcal{A}_{plan}$, it naturally provides relevant sample plans as references for formulating multiple execution plans for the original task.
+We therefore prompt the generation of $N_{p}$ relevant plans independently, yielding $N_{p}$ parallel paths per question.
+This establishes a basis for scaling subsequent agents within document processing workflows, increasing the probability that at least one reasoning path will produce an accurate answer aligned with document semantics.
+(2) [Uncaptioned image] Execution agent: Given that the execution process is divided by step in $\mathcal{A}_{exe}$, we treat each step as an evaluation node.
 For each node, the agent produces $\mathit{N_{e}}$ candidate executions, which are scored using a pretrained reward model.
 The top-scoring candidate is selected as the base node for subsequent steps, with all others rejected.
-(3) Judgment Agent: To judge the correctness of the execution plans and processes, $\mathcal{A}_{judg}$ requires logical analysis and reasoning to comprehensively and accurately detect mistakes.
-We therefore adopt the budget forcing scaling method (Muennighoff et al.
-2025) for this agent, which enforces a minimum number of thinking tokens.
+(3) [Uncaptioned image] Judgment agent: To judge the correctness of the execution plans and processes, $\mathcal{A}_{judg}$ requires logical analysis and reasoning to comprehensively and accurately detect mistakes to avoid the error snowballing in document contexts.
+We therefore adopt the budget forcing scaling method [muennighoff2025s1] for this agent, which enforces a minimum number of thinking tokens.
 When token usage falls below the budget, the agent is encouraged to generate additional thinking tokens, thereby promoting accurate judgments.
-For the answer agent, whose key function is to summarize prior information and generate the final answer, test-time scaling yields only limited improvements, so we do not apply scaling to it.
-Document Non-Document Average
+For the answer agent, whose core function lies in synthesizing prior information and generating the final response, test-time scaling confers merely marginal improvements.
+2.3 Mixed Reward Modeling
+According to the latest research [setlur2025scaling, zhang2025survey], adding RL strategy, e.g., reward modeling, to optimize test-time scaling is significantly superior to fine-tuning or training-free methodologies.
+To optimize our proposed procedural scaling by elevating task-agent-centric and collaborative performance, we design a mixed reward strategy that synthesizes agent-specific rewards and global outcome-driven reward signals.
+Given that the four agents in the framework have distinct functions and reward signal preferences, we first incorporate agent-tailored rewards:
+$r_{plan/exe}$ $=\mathcal{R}_{prm}\left({s}_{i}/{e}_{i}\mid\mathcal{Q},\mathcal{D},\mathcal{P}/\mathcal{E}\right),$ (6)
+$r_{judg/ans}$ $=\mathcal{R}_{orm}\left(\mathcal{J}/\mathcal{O}\mid\mathcal{Q}\right),$ (7)
+where are $\mathcal{R}_{prm}$ and $\mathcal{R}_{orm}$ are multi-modal process reward and outcome reward models, respectively.
+The former yields step-wise process rewards, providing instant feedback for each step and hierarchical rewards for $\mathcal{A}_{plan}$ and $\mathcal{A}_{exe}$, and the latter generates single reward signal for each output of $\mathcal{A}_{judg}$ and $\mathcal{A}_{ans}$.
+Furthermore, we apply a global reward based on the final selected path of the four agents:
+$r_{global}=\mathcal{R}_{orm}\left(\left\{\mathcal{P},\mathcal{E},\mathcal{J},\mathcal{O}\right\}\mid\mathcal{Q},\mathcal{D}\right),$ (8)
+which reinforces incentives for correct paths and moderately mitigates agent self-interest.
+Since individual agents tend to optimize their own local rewards, the integration of a global objective effectively alleviates such self-serving tendencies.
+Please refer to Appendix 7.3 for details.
+Table 1: Comparisons of our MACT and other counterparts on 15 benchmarks.
+The table cells with green, blue, and gray colors denote the best, the second best, and the third best values.
+Results with - indicate exceeding the maximum context token limit or not being equipped with a specific ability.
+$\triangle,\Box,\Diamond$ correspond to the base models of three variants of MACT for comparisons.
+Document Non-Document Avg.
 Text Webpage Chart Table General Mathematical
 DVQA DUDE SVQA MMLong VisMRC InfVQA ChartQA CharXiv TabVQA TabBen SciQA RealQA MVista MVision MVerse
-Average Images Per Question 1.0 5.7 20.0 47.5 1.0 1.2 1.0 1.0 1.0 1.0 0.5 1.0 1.0 1.0 1.0 5.7
-Average Question Length 8.3 8.7 9.0 16.4 10.6 11.5 7.2 20.4 10.5 20.3 12.1 11.1 15.6 42.3 35.7 16.0
-Average Answer Length 2.1 3.4 2.4 2.8 9.5 1.6 1.2 2.9 4.7 8.5 4.4 1.2 1.2 2.2 1.4 3.3
 Generalist
 Closed-Source
-GPT-4o-latest 93.1 52.7 81.0 40.3 86.4 79.2 86.5 78.7 64.2 51.9 83.3 76.2 62.4 30.7 39.4 67.2
-Claude-3.7-Sonnet 94.0 58.1 83.6 33.9 82.5 75.3 92.2 83.5 70.3 54.7 80.4 71.9 69.7 41.2 45.0 69.1
-Gemini-2.0-Pro 91.8 54.3 78.9 32.2 91.4 81.6 88.8 83.1 71.2 59.9 80.9 70.5 74.8 54.2 56.6 71.3
+[…]
+[…]
+[…]
 Size << 20B
@@
 Qwen2.5-VL-7B-Instruct 94.6 62.3 76.2 22.4 86.8 82.8 87.5 68.2 60.4 49.2 74.0 68.4 67.8 25.5 40.8 64.5
-MiMo-VL-7B-SFT 92.9 62.3 74.2 27.5 87.3 82.7 86.1 72.0 57.5 52.5 75.1 67.9 70.7 46.9 53.3 67.3
+[…]
 InternVL3-8B-Instruct 91.1 60.1 72.9 28.6 84.4 78.5 85.4 67.1 59.7 52.4 76.4 66.7 65.5 28.8 38.9 63.8
@@
 LLaVA-OneVision-72B-si 91.7 62.7 63.9 10.6 81.4 75.2 86.1 63.8 52.8 41.5 75.1 70.5 72.5 25.2 27.0 60.0
-Qwen2.5-VL-72B-Instruct 95.7 67.0 82.4 24.9 90.3 87.5 89.5 80.6 68.4 58.1 74.4 74.8 77.6 39.8 46.2 70.5
-InternVL3-78B-Instruct 95.3 68.9 80.7 32.4 90.7 86.2 89.6 77.9 65.5 57.9 78.0 77.9 79.3 44.3 48.7 71.6
+[…]
+[…]
 Llama-3.2-90B-Vision-Instruct 91.4 58.6 76.7 19.5 86.0 82.3 86.3 74.4 60.1 49.4 78.5 68.8 58.7 23.4 35.0 63.3
@@
 Ours
-MACT 96.6 72.5 85.3 43.7 92.0 89.4 91.9 85.2 74.0 57.2 78.5 77.7 81.2 41.8 54.7 74.8
--Qwen2.5-VL-Series-24B +2.0 +10.2 +9.1 +21.3 +5.2 +6.6 +4.4 +17.0 +13.6 +8.0 +4.5 +9.3 +13.4 +16.3 +13.9 +10.3
-MACT 94.4 70.8 83.9 47.4 93.8 88.6 91.4 87.2 71.6 62.7 79.2 76.1 85.4 60.1 65.3 77.2
--MiMo-VL-Series-28B +1.5 +8.5 +9.7 +19.9 +6.5 +5.9 +5.3 +15.2 +14.1 +10.2 +4.1 +8.2 +14.7 +13.2 +12.0 +9.9
-MACT 96.1 69.8 81.3 46.5 91.0 87.7 90.6 85.0 75.3 61.0 81.3 80.1 83.8 45.8 54.0 75.3
--InternVL3-Series-28B +5.0 +9.7 +8.4 +17.9 +6.6 +9.2 +5.2 +17.9 +15.6 +8.6 +4.9 +13.4 +18.3 +17.0 +15.1 +11.5
-Table 1: DVQA, SVQA, MMLong, InfQA, VisMRC, TabVQA, TabBen, SciQA, RealQA, MVista, MVision, and MVerse denote DocVQA, SlideVQA, MMLongBench-Doc, InfographicVQA, VisualMRC, TableVQA-Bench, TableBench, ScienceQA, RealWorldQA, MathVista, Math-Vision, and MathVerse benchmarks, respectively.
-The values with green, blue, and gray colors denote the best, the second best, and the third best values.
-Results with - indicate exceeding the maximum context token limit or not being equipped with a specific ability.
-$\triangle,\Box,\Diamond$ correspond to the base models of MACT for comparisons.
-Experimental Settings
-Training Pipeline
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+3 Experiments
+More details and descriptions about the training pipelines, datasets, evaluations and prompts for each agent are provided in Appendix 8.1, 8.2, 8.3, and 8.4, respectively.
+3.1 Training Pipeline
 Since the functions and optimization goals of each agent are independent and diverse, we design a two-stage SFT and RL pipeline for MACT.
 To understand visual inputs and better accomplish their labors, we choose VLMs for the $\mathcal{A}_{plan}$ and $\mathcal{A}_{exe}$, while LLMs for the $\mathcal{A}_{judg}$ and $\mathcal{A}_{ans}$.
-For all the four agents, we initiate with pretrained models, and we select three groups of small-parameter base models for different variants of MACT: (1) Qwen2.5-VL Series Based (Bai et al.
-2025; Yang et al.
-2025): Qwen2.5-VL-7B-Instruct and Qwen2.5-7B/3B-Instruct; (2) MiMo-VL Series Based (Xiaomi 2025; Xia et al.
-2025): MiMo-VL-7B-SFT and MiMo-7B-SFT; (3) InternVL3 Series Based (Zhu et al.
-2025): InternVL3-9B/8B/2B-Instruct.
+For all the four agents, we initiate with pretrained models, and we select three groups of small-parameter base models for different variants of MACT: (1) Qwen2.5-VL series based [bai2025qwen2, qwen2025qwen25technicalreport]: Qwen2.5-VL-7B-Instruct and Qwen2.5-7B/3B-Instruct; (2) MiMo-VL series based [coreteam2025mimovl, xia2025mimo]: MiMo-VL-7B-SFT and MiMo-7B-SFT; (3) InternVL3 series based [zhu2025internvl3]: InternVL3-9B/8B/2B-Instruct.
 In the first SFT stage, we initially train a tuned 11B/7B/7B VLM on the selected document-based or non-document-based datasets, mixing data with or without CoT.
 It aims to enhance their visual understanding and reasoning abilities, providing more robust models for future multi-agent collaboration.
-Next, we fine-tune an 8B/7B/7B LLM on judgment labels generated via GPT-4o (Hurst et al.
-2024) and rule-based verifications.
+Next, we fine-tune an 8B/7B/7B LLM on judgment labels generated via GPT-4o [hurst2024gpt] and rule-based verifications.
 Finally, we fine-tune another 3B/3B/7B LLM on the outputs from preceding agents and ground-truths to enhance its summary ability.
 The fine-tuned VLM are applied as $\mathcal{A}_{plan}$ and $\mathcal{A}_{exe}$ respectively in the subsequent stage, while the two LLMs function as the $\mathcal{A}_{judg}$ and $\mathcal{A}_{ans}$.
-In the second RL stage, we generate reward signals based on pretrained reward models and optimize our model via GRPO (Shao et al.
-2024).
-Process reward model VisualPRM (Wang et al.
-2025c) is utilized for the step-by-step reward signals of $\mathcal{A}_{plan}$ and $\mathcal{A}_{exe}$, while Skywork-VL-Reward (Wang et al.
-2025d) is used for the rewards generation of $\mathcal{A}_{judg}$ and $\mathcal{A}_{ans}$.
-More training details are in the Appendix.
-Datasets
-To train our model and comprehensively, objectively assess its visual capabilities, we selected 15 datasets spanning four real-world document types and two non-document types.
-The four document types are comprised of: (1) Text-Based: DocVQA (Mathew, Karatzas, and Jawahar 2021), DUDE (Van Landeghem et al.
-2023), SlideVQA (Tanaka et al.
-2023), MMLongBench-Doc (Ma et al.
-2025); (2) Webpage-Based: VisualMRC (Tanaka, Nishida, and Yoshida 2021), InfographicVQA (Mathew et al.
-2022); (3) Chart-Based: ChartQA (Masry et al.
-2022), CharXiv (Wang et al.
-2024d); (4) Table-Based: TableVQA-Bench (Kim, Yim, and Song 2024), and TableBench (Wu et al.
-2025).
-To retain generalization and enhance reasoning, two non-document types are involved: (1) General: ScienceQA (Saikh et al.
-2022), RealWorldQA (xAI 2024); (2) Mathematical: MathVista (Lu et al.
-2024a), Math-Vision (Wang et al.
-2024b), and MathVerse (Zhang et al.
-2024c).
-We adhere to the original training and testing splits and use the test splits as evaluation benchmarks.
-Each instance in these datasets consists of a visual input and a question (except ScienceQA), spanning difficulty levels from easy to hard and encompassing various question types.
-More details are provided in the Appendix.
-Evaluations
-Following (Li et al.
-2025b), we use the reported initial results of reproducible models where configurations align with ours.
-For other models, we utilize LMMs-Eval (Zhang et al.
-2024b) for fair comparisons on both natively supported and our registered benchmarks.
-For most benchmarks, we employ GPT-4o (Hurst et al.
-2024) as a judge model to evaluate the correctness of the generated answers, which is achieved in LMMs-Eval.
-For the remaining benchmarks, we utilize their original evaluation metrics, e.g., ANLS and F1.
-More details can be seen in the Appendix.
-Results and Discussions
-Main Results
-We evaluate MACT across 15 benchmarks: 10 document-based benchmarks to assess its document understanding performance; 2 general benchmarks to gauge generalization ability, mitigating overfitting to specific scenarios; and 3 mathematical benchmarks to evaluate complex reasoning capabilities.
-For comparison, we select state-of-the-art (SOTA) methods—encompassing both generalist models and document-specific specialists—and categorize them by parameter size.
-As shown in Table 1, the MACT-MiMo-VL-Series-28B variant delivers the best average performance, followed by MACT-InternVL3-Series-28B and MACT-Qwen2.5-VL-Series-24B, which rank second and third, respectively.
+In the second RL stage, we generate reward signals based on pretrained reward models and optimize our model via GRPO [shao2024deepseekmath].
+Process reward model VisualPRM [wang2025visualprm] is utilized for the step-by-step reward signals of $\mathcal{A}_{plan}$ and $\mathcal{A}_{exe}$, while Skywork-VL-Reward [wang2025skywork] is used for the rewards generation of $\mathcal{A}_{judg}$ and $\mathcal{A}_{ans}$.
+3.2 Benchmarks
+To comprehensively and objectively evaluate our model’s document-based capabilities, we selected 15 datasets encompassing four real-world document categories and two non-document general categories.
+The four document types are comprised of: (1) Text-based: DocVQA [mathew2021docvqa], DUDE [van2023document], SlideVQA [tanaka2023slidevqa], MMLongBench-Doc [ma2025mmlongbench]; (2) Webpage-based: VisualMRC [tanaka2021visualmrc], InfographicVQA [mathew2022infographicvqa]; (3) Chart-based: ChartQA [masry2022chartqa], CharXiv [wang2024charxiv]; (4) Table-based: TableVQA-Bench [kim2024tablevqa], and TableBench [wu2025tablebench].
+To validate our document-centric paradigm without sacrificing capabilities in general domains, two non-document types are involved: (1) General: ScienceQA [saikh2022scienceqa], RealWorldQA [2024XAIrealworld]; (2) Mathematical: MathVista [lu2024mathvista], Math-Vision [wang2024measuring], and MathVerse [zhang2024mathverse].
+We adhere to the original training and testing splits and use the testing splits as evaluation benchmarks.
+Each instance in these datasets consists of a visual input sequence and a question, spanning difficulty levels from easy to hard and encompassing various question types for comprehensive evaluation.
+3.3 Evaluations
+For other reproducible models, we utilize LMMs-Eval [zhang2024lmms] for fair comparisons on both natively supported and our registered benchmarks.
+For most benchmarks, we employ GPT-4o [hurst2024gpt] as a judge model to evaluate the correctness of the generated answers based on LMMs-Eval.
+For the remaining benchmarks, we utilize their original evaluation metrics, e.g., ANLS and F1, as detailed in Tab.
+8.
+4 Results and Discussions
+4.1 Main Results
+We evaluate MACT across 15 benchmarks: 10 document-based benchmarks to assess its document understanding performance; 2 general benchmarks and 3 mathematical benchmarks are designed to verify that it retains full capabilities in non-document-centric and reasoning task settings.
+For comparison, we select state-of-the-art (SOTA) methods, encompassing both generalist and specialist models, and categorize them by parameter size.
+Benchmark Results.
+As shown in Tab.
+1, the MACT-MiMo-VL-Series-28B variant delivers the best average performance, followed by MACT-InternVL3-Series-28B and MACT-Qwen2.5-VL-Series-24B, which rank second and third, respectively.
 Despite having fewer than 30B parameters, MACT models outperform all comparative methods with under 100B parameters, as well as closed-source models.
@@
 Notably, in MMLongBench-Doc, which features the longest visual context, and across the three mathematical reasoning benchmarks, MACT-MiMo-VL-Series-28B outperforms the second-highest scorer by 7.1%, 10.6%, 5.9%, and 8.7%, respectively.
-These results highlight the superior overall performance of MACT, as well as its strong capabilities in long-context understanding and reasoning.
-Radar charts in Figure 1 compare three variants of MACT with their base models and larger models from the same series.
-As shown, MACT models significantly outperform their base models by 10.3%, 9.9%, and 11.5% on average across the 15 benchmarks.
-Compared to their corresponding larger-parameter counterparts, i.e., Qwen2.5-VL-72B-Instruct and InternVL3-78B-Instruct, our MACT-Qwen2.5-VL-Series-24B and MACT-InternVL3-Series-28B achieve average performance gains of at least 6.6% and 3.7%, respectively, with more pronounced improvements observed on long-context and reasoning tasks.
-Additional Quantitative Analysis
-To further demonstrate the effectiveness of our methodologies, we perform ablation experiments and analyses on the proposed multi-agent collaboration, reward modeling, and test-time scaling strategies.
-Experiments here are conducted using MACT-Qwen2.5-VL-Series-24B.
-Specifically, we select three lowest-scoring benchmarks, which are characterized by long visual contexts and reasoning demands, and the average results across all benchmarks to assess their impact on both specific abilities and overall performance.
-Ablation Studies
-MMLong TabBen MVision Average
-w/o Multi-Agent Collaboration 32.5 $\downarrow$11.2 50.8 $\downarrow$  6.4 32.4 $\downarrow$  9.4 66.2 $\downarrow$  8.6
-w/o Multi-Agent Collaboration 24.7 $\downarrow$19.0 44.4 $\downarrow$12.8 26.0 $\downarrow$15.8 58.6 $\downarrow$16.2
-w/o Mixed Reward Modeling 38.3 $\downarrow$  5.4 54.2 $\downarrow$  3.0 36.7 $\downarrow$  5.1 71.4 $\downarrow$  3.4
-w/o Agent-Wise Hybrid Scaling 34.9 $\downarrow$  8.8 50.8 $\downarrow$  6.4 34.5 $\downarrow$  7.3 71.1 $\downarrow$  3.7
+These results highlight that document-based tasks, far from being monolithic constructs, are better suited to procedural scaling than to monolithic models with larger parameters, supporting flexible adaptation to functional entities and computational resource allocation.
+Comparisons with Base Models.
+As shown in Fig.
+1, radar charts compare three variants of our proposed MACT framework with their base models and larger models from the same series.
+It is observed that MACT models significantly outperform their base models by 10.3%, 9.9%, and 11.5% on average across the 15 benchmarks.
+Besides, compared to their corresponding larger monolithic models, i.e., Qwen2.5-VL-72B-Instruct and InternVL3-78B-Instruct, our MACT-Qwen2.5-VL-Series-24B and MACT-InternVL3-Series-28B achieve average performance gains of at least 6.6% and 3.7%, respectively, with more pronounced improvements observed on long-context and reasoning tasks.
+These results demonstrate that our method, equipped with procedural scaling paradigm, is superior to monolithic paradigm in the document-based tasks and also more general domains.
+4.2 Additional Quantitative Analysis
+To further demonstrate the effectiveness of our proposed procedural scaling paradigm, we perform ablation experiments and analyses of scaling components.
+Experiments here are conducted using MACT-Qwen2.5-VL-Series-24B, and the training pipelines and training data are consistent with the main experiment.
+We select three lowest-performing benchmarks, distinguished by long visual contexts and reasoning-intensive demands, and the average performance across all benchmarks, aiming to quantify their impacts on both specialized capabilities.
+Table 2: Results of ablations on multi-agent collaboration, agent-wise adaptive test-time scaling, and mixed reward modeling.
+MMLong TabBen MVision Avg.
+Monolithic 32.5 50.8 32.4 66.2
+w/o Multi-Agent Collaboration 24.7 44.4 26.0 58.6
+w/o Agent-Wise Adaptive Scaling 34.9 50.8 34.5 71.1
+w/o Mixed Reward Modeling 38.3 54.2 36.7 71.4
 MACT 43.7 57.2 41.8 74.8
-Table 2: Results of ablations. $\dagger$ denotes a single agent that directly outputs answers, while $\ddagger$ indicates a single agent that incorporates all prompts as a workflow.
-$\mathcal{A}_{plan}$ + $\mathcal{A}_{exe}$ $\mathcal{A}_{judg}$ $\mathcal{A}_{ans}$ MMLong TabBen MVision Average
+Table 3: Results of different combinations of agents.
+$\mathcal{A}_{plan}$ + $\mathcal{A}_{exe}$ $\mathcal{A}_{judg}$ $\mathcal{A}_{ans}$ MMLong TabBen MVision Avg.
 36.3 48.5 34.9 68.4
@@
 43.7 57.2 41.8 74.8
-Table 3: Results of different combinations of agents.
-The ablation settings are as follows: (1) w/o Multi-Agent Collaboration: Use a single agent to directly execute tasks and output answers, with our proposed scaling and reward strategies retained.
+Ablation Studies.
+The ablation settings are as follows: (1) Monolithic: Use a monolithic model to directly execute tasks and output answers, with particular scaling and reward modeling retained.
 (2) w/o multi-agent collaboration: Use a single agent, which incorporates all prompts from our four proposed agents as a unified workflow, with our proposed scaling and reward strategies retained.
-(3) w/o Agent-Wise Hybrid Scaling: Retain the multi-agent collaborative framework and apply the mixed reward strategy directly, without test-time scaling.
-(4) w/o Mixed Reward Modeling: Retain the multi-agent collaborative framework and agent-wise hybrid scaling, without reward signals.
-As shown in Table 2, the multi-agent collaborative approach contributes the most, outperforming the single-agent system by 8.6%.
-Notably, consolidating all functions into a single model results in suboptimal performance, which is even worse than the base model, highlighting the necessity of a multi-agent collaborative framework.
-Both the mixed reward modeling and agent-wise hybrid scaling strategies boost average accuracy by at least 3%.
-Specifically, the latter yields more significant gains on complex tasks compared to average ones, with especially notable improvements in reasoning tasks.
-Analysis of Multi-Agent Collaboration
-To assess the utility of each agent and their collaboration, we perform additional experiments on different agent combinations, which are listed in Table 3.
-The combination of $\mathcal{A}_{plan}$ and $\mathcal{A}_{exe}$ results in a 3.9% average increase over the base model, while adding $\mathcal{A}_{judg}$ further boosts performance by 5.5%.
-$\mathcal{A}_{ans}$ extraly improves the scores by 0.9%, with this gain expanding to 1.5% for long visual context.
-Additionally, as in Figure 4(a), we compare our judgment agent strategy with the other two counterparts.
-Our proposed strategy outperforms the others by at least 2.6% on average, while requiring an average of 0.3 fewer corrections.
-Moreover, our method achieves optimal correction results when the maximum number of corrections $N_{c}$ is set to 3, whereas the alternatives require $N_{c}$ to be set to 5, highlighting its greater accuracy and efficiency.
-It could also be observed that when setting a higher $N_{c}$, the performance will not improve continuously, primarily because excessive corrections may confuse the agent and obscure correct answers.
-MMLong TabBen MVision Average
-No Reward 38.3 54.2 36.7 71.4
-Agent-Specific Reward 42.8 56.3 40.5 72.7
-Global Reward 34.1 51.6 35.1 70.2
+(3) w/o agent-wise adaptive test-time scaling: Retain the multi-agent collaborative framework and apply the mixed reward strategy directly, without test-time scaling.
+(4) w/o mixed reward modeling: Retain the multi-agent collaborative framework and agent-wise adaptive test-time scaling, without reward signals.
+As reported in Tab.
+2, our document-centric procedural scaling paradigm, outperforming the monolithic system by 8.6%.
+Notably, the integration of all functionalities into a single model results in suboptimal performance, even outperforming the base model negatively, which indicates the inherent limitation of monolithic paradigms in document understanding and reasoning.
+Both the agent-wise adaptive test-time scaling strategy and mixed reward modeling contribute to an average accuracy increase of at least 3%.
+Specifically, the former yields more remarkable gains on reasoning tasks in documents than average ones, with particularly striking improvements that underscore how adaptive scaling reduces cognitive overload.
+Table 4: Results of different setting of test-time scaling strategies.
+MMLong TabBen MVision Avg.
 No Scaling 34.9 50.8 34.5 71.1
@@
 Internal Scaling 41.8 55.9 41.6 72.3
-Agent-Wise Hybrid 43.7 57.2 41.8 74.8
-Scaling & Mixed Reward
-Table 4: Results of different reward modeling and test-time scaling strategies.
+Agent-Wise Adaptive 43.7 57.2 41.8 74.8
+Table 5: Results of different setting of reward modeling.
+MMLong TabBen MVision Avg.
+No Reward 38.3 54.2 36.7 71.4
+Agent-Specific Reward 42.8 56.3 40.5 72.7
+Global Reward 34.1 51.6 35.1 70.2
+Mixed Reward 43.7 57.2 41.8 74.8
+Analysis of Multi-Agent Collaboration.
+To assess the utility of each agent and the effectiveness their procedural collaboration, we perform additional experiments on different agent combinations, which are demonstrated in Tab.
+3.
+The combination of $\mathcal{A}_{plan}$ and $\mathcal{A}_{exe}$, as a basic procedural multi-agent system, results in a 3.9% average increase over the base model, while adding $\mathcal{A}_{judg}$ further boosts performance by 5.5%.
+$\mathcal{A}_{ans}$ additionally improves the scores by 0.9% on the average.
+Additionally, as in Fig.
+4(a), we compare our judgment agent strategy with the other two counterparts, i.e., internal correction and agent for both judgment and correction.
+Our proposed strategy outperforms the others by at least 2.6% on average, while requiring an average of 0.3 fewer corrections.
+Moreover, our method achieves optimal correction results when the maximum number of corrections $N_{c}$ is set to 3, whereas the alternatives require $N_{c}$ to be set to 5, highlighting its greater accuracy and efficiency.
+It could also be observed that when setting a higher $N_{c}$, the performance will not improve continuously, primarily because excessive corrections may confuse the agent and obscure correct answers.
 (a) Analysis of Corrections
-(b) Impact of $N_{p}$ and $N_{e}$
-Figure 4: (a) Line graph shows the impact of various maximum numbers of corrections, with dashed and dotted lines denoting average values across all and three selected benchmarks, respectively.
+(b) Impact of $\mathit{N_{p}}$ and $\mathit{N_{e}}$
+Figure 4: (a) Line graph shows the impact of various maximum numbers of corrections, with solid and dashed lines denoting average values across all and three selected benchmarks, respectively.
 The bar charts show the average judgment numbers when the maximum is set to 3.
-(b) The line graphs represent the impacts of $N_{p}$ and $N_{e}$.
-Analysis of Reward modeling and Test-Time Scaling
-We perform experiments on our proposed mixed reward modeling and agent-wise hybrid test-time scaling.
-For the former, we separately use the agent-wise reward and global reward for comparison.
-For the latter, we benchmark our method with the other four effective test-time scaling strategies.
-Here, we select budget forcing (Muennighoff et al.
-2025) as the internal scaling.
+(b) The line graphs represent the impacts of the number of generated plans $\mathit{N_{p}}$ and candidate executions $\mathit{N_{e}}$.
+Analysis of Agent-Wise Adaptive Test-Time Scaling and Mixed Reward modeling.
+We perform experiments on our proposed agent-wise adaptive test-time scaling and mixed reward modeling.
+For the former, we benchmark our method with the other four effective test-time scaling strategies, as listed in Tab.
+4.
+Here, we select budget forcing [muennighoff2025s1] as the internal scaling.
+Our agent-wise adaptive scaling strategy based on the computational resources and cognitive load of each agent, outperforms all four existing strategies and improves the average scores by at least 1.8%.
+As shown in Fig.
+4(b), higher values of $\mathit{N_{p}}$ and $\mathit{N_{e}}$ increase the likelihood of finding the correct answers, which leads to more attempts at plans and execution steps.
+For the latter, we separately use the agent-wise reward and global reward for comparison, as reported in Tab.
+5.
 Although the improvement from global reward is relatively limited, it avoids the selfishness of agents that only use agent-specific reward, yielding a further 1.3% increase in average scores.
-Our strategy, which designs tailored scaling forms for each agent according to its functions and specificity, outperforms all four existing strategies and improves the average scores by at least 1.8%.
-As demonstrated in Figure 4(b), higher values of $N_{p}$ and $N_{e}$ increase the likelihood of finding the correct answers, which leads to more attempts at plans and execution steps.
-Conclusion
-Our proposed MACT is an innovative multi-agent collaboration framework with a test-time scaling strategy for visual document understanding and question answering.
-Comprising four collaborative agents, this framework achieves effective functional role division through synergistic collaboration.
+5 Conclusion
+Our proposed MACT is an innovative multi-agent collaboration framework with adaptive test-time scaling strategy for visual document understanding and reasoning, shifting the monolithic scaling into procedural scaling.
+It comprises four collaborative agents that deconstructs the workflow of document analysis, achieving effective role division and collaboration.
 Notably, the judgment agent strategy outperforms existing self-correction mechanisms while requiring fewer corrections.
-Additionally, mixed reward modeling and agent-wise hybrid scaling are particularly well-suited for multi-agent systems, as they strike a balance between agent-specific capabilities and collaborative efficiency.
-Extensive comparative experiments and supplementary analysis validate the superiority and effectiveness of our MACT.
-Furthermore, this work constitutes a meaningful exploration of multi-agent frameworks and test-time scaling strategies, underscoring promising directions for future research.
-Appendix
-Appendix A Related Works
-General VLMs
-Recent advancements in general VLMs have largely bridged the gap between visual and linguistic information, demonstrating impressive abilities in visual understanding, reasoning, and question answering.
-For example, latest GPT-4o (Hurst et al.
-2024), GPT-4V (OpenAI 2024), Gemini-2.0-Pro (DeepMind 2025), DeepSeek-VL2 (Wu et al.
-2024), Qwen2.5-VL (Bai et al.
-2025), Claude-3.7-Sonnet (Anthropic 2024), InternVL-3 (Zhu et al.
-2025), Llama-3.2-Vision (Grattafiori et al.
-2024), Ovis2 (Lu et al.
-2024b) and MiMo-VL (Xiaomi 2025) contribute to push the performance boundaries in many aspects.
-However, there is still room for improvement for these general VLMs in terms of specific downstream tasks in various scenarios.
-Furthermore, their derivative models, such as LLaVA-OneVision (Li et al.
-2025b), Monkey (Li et al.
-2024), Emu3 (Wang et al.
-2024c), InternLM-XComposer2-4KHD (Dong et al.
-2024a) and mPLUG-Owl3 (Ye et al.
-2025), contribute to push the performance boundaries in many aspects.
-Specialized VLMs
-To better promote visual understanding and question answering, especially in document-based tasks, such as texts, webpages, charts, and tables, researchers have designed numerous task-specialized VLMs.
-For instance, TGDoc (Wang et al.
-2023) and MMCA (Liu et al.
-2024a) are explorations of document-based or chart-based understanding via large-scale collected instruction tuning data, and DocPedia (Feng et al.
-2024) deals with the visual inputs at the frequency domain.
-To improve the comprehension of screenshots and webpages, UReader (Ye et al.
-2023) designs a cropping strategy, while CogAgent (Hong et al.
-2024) and TextMonkey (Liu et al.
-2024b) use two resolution branches or a shifted window to accommodate document-oriented tasks.
-The series of works by mPLUG-DocOwl (Hu et al.
-2024a, b) design a high-resolution compressor for multi-page document understanding with reduced visual token costs.
-Besides, ColPali (Faysse et al.
-2025), M3DocRAG (Cho et al.
-2024), and MDocAgent (Han et al.
-2025) introduce efficient and novel retrieval mechanisms for VLVMs to comprehend long documents.
-Multi-Agent Models
-Agentic large models have gained tremendous popularity; however, single-agent systems are increasingly unable to meet the demands of complex and novel tasks.
-Thus, communicative agents’ mechanisms are implemented to improve cooperation and reaction between agents by two primary forms, i.e., debate (Du et al.
-2024; Zhang et al.
-2024a; Yue et al.
-2025; Pan et al.
-2024) and competition (Cheng et al.
-2024; Chen et al.
-2024a).
-Although the domain of LLM-based multi-agent models has expanded to include code generation (Islam, Ali, and Parvez 2024; Tao et al.
-2024), financial decision-making (Yu et al.
-2024b), and other areas, they still suffer from a lack of visual understanding.
-Thus, more recently, some recent works attempt to explore how multi-agent interaction benefits VLMs.
-For instance, Insight-V (Dong et al.
-2024b) divides the functions on reasoning and summary agents, while Mobile-Agent-v2 (Wang et al.
-2024a) and  (Li et al.
-2025c) introduce specialized agents for planning.
-To accommodate different tasks, METAL (Li et al.
-2025a), WebPilot (Zhang et al.
-2025b), ViDoRAG (Wang et al.
-2025b), MatPlotAgent (Yang et al.
-2024), and VipAct (Zhang et al.
-2024d) further break down the labor of a single agent into more manageable sub-instances and specialise the function of each role-tailored agent.
-Additionally, MuMA-ToM (Shi et al.
-2025) and MDocAgent (Han et al.
-2025) design separate image and text agents and then fuse them, ignoring the correlations between the two modalities.
-However, a practical multi-agent collaboration framework for visual document understanding and question answering is still deficient.
-Test-Time Scaling
-As a promising and burgeoning research focus, test-time scaling can further elicit various capabilities of models with relatively lower computational cost (Zhang et al.
-2025a).
-The scaling could be categorized into four types: parallel scaling (Renze 2024; Brown et al.
-2024; Wang et al.
-2025a; Liu et al.
-2025), generating multiple candidates and aggregating them; sequential scaling (Gou et al.
-2024; Chen et al.
-2024b; Yu et al.
-2024a), iteratively updating intermediate states and triggering self-correction; Hybrid scaling (Besta et al.
-2024; Guan et al.
-2025; Snell et al.
-2025) is another realization, which exploits their complementary benefits; internal scaling, which autonomously allocates computational tokens, e.g.
-budget forcing strategy (Muennighoff et al.
-2025; Son et al.
-2025; Li et al.
-2025a).
-To further enhance the ability of models to scale at test time, supervised fine-tuning (SFT) and reinforcement learning (RL) are also widely employed to improve scaling performance.
-For example, an intuitive and efficient strategy is to set a reward model, such as process reward modeling (Snell et al.
-2025; Zhao et al.
-2025) and outcome reward modeling (Liu et al.
-2025; Zuo et al.
-2025; Setlur et al.
-2025), offering quantitative reward values for test-time scaling.
-Figure 5: Illustrations of representative samples from our selected 15 benchmarks on both document and non-docuement tpyes.
-Appendix B Experimental Settings
-Training Pipeline
-In this work, we design a two-stage training pipeline to harness the full potential of our framework, specifically the SFT stage and RL stage.
-Moreover, we select three groups of base models for different variants, which are shown in Table 5.
-To enhance the general abilities, during the first stage, we employ AdamW optimizer with $\beta=\left(0.9,0.95\right)$, and the initial learning rate and weight decay are set to 2e52e{-5} and $0.1$ respectively.
-We train our model for one epoch with a batch size of 128.
-In the second RL stage, we adopt the GRPO (Shao et al.
-2024) strategy for further optimisation.
-Following its original configurations, we set the learning rate, KL coefficient, and number of iterations as e6e{-6}, $0.04$, and $1$, respectively.
-The reward signals are generated by VisualPRM (Wang et al.
-2025c) and Skywork-VL-Reward (Wang et al.
-2025d); the former provides step-by-step signals for planning and execution agents, while the latter generates rewards for judgment and answer agents.
-During the process, we set the temperature parameter to $0.75$ to ensure diversity and creativity in the generated solutions.
-It should be noted that reward models and our framework are loaded on different servers, thereby avoiding the repeated loading and offloading of model weights.
-All the training and inference are conducted on 8 NVIDIA A100 GPUs.
-Type Qwen2.5-VL Series MiMo-VL Series InternVL3 Series
-(Bai et al. 2025; Yang et al. 2025) (Xiaomi 2025; Xia et al. 2025) (Zhu et al. 2025)
-Planning Agent VLM Qwen2.5-VL-7B-Instruct MiMo-VL-7B-SFT InternVL3-9B-Instruct
-Execution Agent VLM Qwen2.5-VL-7B-Instruct MiMo-VL-7B-SFT InternVL3-9B-Instruct
-Judgment Agent LLM* Qwen2.5-7B-Instruct MiMo-7B-SFT InternVL3-8B-Instruct
-Answer Agent LLM* Qwen2.5-3B-Instruct MiMo-7B-SFT InternVL3-3B-Instruct
-Size - 7B+7B+7B+3B 7B+7B+7B+7B 9B+9B+8B+2B
-=24B =28B =28B
-Table 5: The base models of each variant of MACT. * indicates that we use VLMs, which remain powerful textual ability, because in InternVL3 series the model family does not provide LLMs.
-Datasets
-Our selected datasets cover four document types: text-based, webpage-based, table-based, and chart-based benchmarks, and also two non-document types: general and mathematical benchmarks.
-For each type, two to four datasets are included for both training and evaluation benchmarks.
-To better illustrate the datasets, statistical information provided in Table 6 and a brief introduction of each dataset are listed as follows:
-Dataset Type Num I Num Q Avg Q Avg A Avg I Per Q
-DocVQA (Mathew, Karatzas, and Jawahar 2021) Text 12K 50K 8.3 2.1 1.0
-DUDE (Van Landeghem et al. 2023) Text 5K 41K 8.7 3.4 5.7
-SlideVQA (Tanaka et al. 2023) Text 52K 15K 9.0 2.4 20.0
-MMLongBench-Doc (Ma et al. 2025) Text 6K 1K 16.4 2.8 47.5
-VisualMRC (Tanaka, Nishida, and Yoshida 2021) Webpage 10K 30K 10.6 9.5 1.0
-InfographicVQA (Mathew et al. 2022) Webpage 5K 30K 11.5 1.6 1.2
-ChartQA (Masry et al. 2022) Chart 22K 34K 7.2 1.2 1.0
-CharXiv (Wang et al. 2024d) Chart 2K 13K 20.4 2.9 1.0
-TableVQA-Bench (Kim, Yim, and Song 2024) Table 1K 2K 10.5 4.7 1.0
-TableBench (Wu et al. 2025) Table 21K 4K 20.3 8.5 1.0
-ScienceQA (Saikh et al. 2022) General 10K 21K 12.1 4.4 0.5
-RealWorldQA (xAI 2024) General 1K 1K 11.1 1.2 1.0
-MathVista (Lu et al. 2024a) Mathematical 5K 6K 15.6 1.2 1.0
-Math-Vision (Wang et al. 2024b) Mathematical 3K 3K 42.3 2.2 1.0
-MathVerse (Zhang et al. 2024c) Mathematical 3K 16K 35.7 1.4 1.0
-Table 6: Statistical information of our selected datasets. Here, Num and Avg indicate the total number and the average length, while I, Q, and A represent images, questions, and answers, respectively.
-  • DocVQA (Mathew, Karatzas, and Jawahar 2021).
-It is a classic document-based dataset, with more than 12K images from various documents and 50K related questions.
-Each question corresponds to a single-page document, which requires relatively simple visual understanding.
-The average lengths of questions and answers are 8.3 and 2.1 words.
-  • DUDE (Van Landeghem et al.
-2023).
-It is a multi-page document question answering dataset, comprising 41,541 questions and over 5,000 documents.
-The average number of pages in the document of this dataset is 5.72, while the average lengths of the questions and answers are 8.7 and 3.4, respectively.
-  • SlideVQA (Tanaka et al.
-2023).
-It is a slide question answering dataset, which incorporates 52K images from slides and about 15K questions.
-The average lengths of the questions and answers are 9.0 and 2.4 words, and they include a certain proportion of multi-hop and numerical reasoning.
-  • MMLongBench-Doc (Ma et al.
-2025).
-It introduces a multi-page dataset for document understanding, consisting of about 6K document images and 1K questions.
-Depending on multi-hop evidence, the questions and answers have 16.4 and 2.8 words on average, and almost half of the questions are cross-page, while one-fifth of the questions are unanswerable.
-  • VisualMRC (Tanaka, Nishida, and Yoshida 2021).
-It collects 10K images sourced from 35 domains of webpages and more than 30K annotated questions with an average length of 10.6 words.
-The answers, with an average length of 9.5 words, in this dataset include both word spans in the context and the layouts of the webpages, which are more complex and diverse, with more abstract questions.
-  • InfographicVQA (Mathew et al.
-2022).
-It consists of more than 5K images collected from webpages and 30K questions, constructed with textual, graphical, and visual elements.
-Most questions in this dataset are extractive and numerical, testing basic reasoning and arithmetic skills, with 11.5 and 1.6 words in questions and answers on average.
-  • ChartQA (Masry et al.
-2022).
-It introduces a dataset with 22K images and 34K questions, involving bar, line, pie charts, and composite plots.
-The questions in this dataset require visual, logical, and arithmetic ability, with an average length of questions and answers of 7.2 and 1.2 words.
-  • CharXiv (Wang et al.
-2024d).
-It presents a dataset for chart understanding and answering, with 2K chart images and 13K curated questions.
-It includes both descriptive and reasoning sub-datasets, and we merge them.
-The questions and answers have 22.6 and 2.8 words evenly.
-  • TableBench (Wu et al.
-2025).
-It is a dataset with tabular and textual hybrid contents, constructed with over 4K tables and 21K questions, with numerical reasoning, such as computing, comparison, and sorting.
-Its questions and answers have lengths of 20.3 and 8.5 words on average.
-  • TableVQA-Bench (Kim, Yim, and Song 2024).
-It establishes a dataset for table-based question answering, extended with four existing sources with 1K tables and 2K questions.
-Consisting of diverse types of questions, the average lengths of questions and answers are 10.5 and 4.7 words.
-  • ScienceQA (Saikh et al.
-2022).
-It is a dataset that includes about 21K multiple-choice questions based on more than 10K images from diverse scientific fields, and the length of the questions and answers is 12.1 words on average.
-  • RealWorldQA (xAI 2024). It consists of about 1K questions and 1K images. The average length of the questions in this dataset is 11.1 words, while that of the answers is only 1.2 words.
-  • MathVista (Lu et al.
-2024a).
-It proposed a dataset integrating diverse mathematical and visual tasks with 5K images and 6K questions, demanding fine-grained visual understanding and compositional reasoning.
-On average, the length of the questions and answers is 15.6 and 1.2 words, respectively.
-  • Math-Vision (Wang et al.
-2024b).
-It is a dataset integrating diverse mathematical tasks with about 3K images and 3K questions from real exams, spanning 16 distinct mathematical disciplines and graded across 5 levels of difficulty.
-It has the longest average length of questions (42.3 words) in our selected datasets, and 2.2 words in answers on average.
-  • MathVerse (Laurençon et al.
-2024).
-It is a mathematical dataset with about 3K images and 16K questions.
-On average, the lengths of the questions and answers in this dataset are 35.7 and 1.4 words, respectively.
-For some datasets, we use pre-treatments to handle the visual inputs.
-For the document in some datasets, which is provided in PDF format, we transform these files into visual images as input.
-Besides, we design a standardised visualisation process for part of the tables in the TableBench (Wu et al.
-2025) dataset, which only contains HTML format data without authentic visual images.
-Additionally, we utilize the PDF2Image library to convert PDF files into images for input.
-Evaluations
-We onboard the unavailable models in LMMs-Eval (Zhang et al.
-2024b) to evaluate their performances with the consistent configurations.
-We directly use the natively supported benchmarks in the LMMs-Eval framework, namely DocVQA (Mathew, Karatzas, and Jawahar 2021), InfographicVQA (Mathew et al.
-2022), ChartQA (Masry et al.
-2022), ScienceQA (Saikh et al.
-2022), RealWorldQA (xAI 2024), MathVista (Lu et al.
-2024a), and MathVerse (Zhang et al.
-2024c).
-In terms of unavailable benchmarks in LMMs-Eval, namely DUDE (Van Landeghem et al.
-2023), SlideVQA (Tanaka et al.
-2023), MMLongBench-Doc (Ma et al.
-2025), VisualMRC (Tanaka, Nishida, and Yoshida 2021), CharXiv (Wang et al.
-2024d), TableVQA-Bench (Kim, Yim, and Song 2024), TableBench (Wu et al.
-2025), and Math-Vision (Wang et al.
-2024b), we onboard the benchmarks and register them by YAML configurations.
-To evaluate the performance of models on these benchmarks fairly and equitably, we employ a range of evaluation metrics.
-For DocVQA (Mathew, Karatzas, and Jawahar 2021), DUDE (Van Landeghem et al.
-2023), InfographicVQA (Mathew et al.
-2022) benchmarks, we follow the original evaluation metrics, i.e., ANLS (Average Normalised Levenshtein Similarity) (Biten et al.
-2019), to measure the difference between the predicted answers and ground truths.
-Moreover, the F1 score is utilized for SlideVQA (Tanaka et al.
-2023).
-For other selected benchmarks, we use powerful GPT-4o (Hurst et al.
-2024) as a judge model to evaluate the answers by LMMs-Eval, which generates scores from 1 to 10, and then the scores will be scaled to an accuracy between 0 and 100 in the dimension of benchmarks.
-Figure 6: Prompts for the four agents.
-Prompt Templates
-As shown in Figure 6, there are prompts for the four agents, including the system prompts and the prompts for each process.
-Apart from the system prompts, $\mathcal{A}_{plan}$ and $\mathcal{A}_{exe}$ have two extra prompts to accomplish their functions, while $\mathcal{A}_{judg}$ and $\mathcal{A}_{ans}$ have only one.
-In the separating prompt of $\mathcal{A}_{exe}$, each execution unit is comprised of: (1) a specific definition; (2) expected target and output; (3) existing inputs or results from the previous step.
-Appendix C Results and Discussions
-Additional Quantitative Results
-For the line graphs and bar charts shown in Figure 4(a) and the line graphs in Figure 4(b), we list the quantitative results in Table 7 and Table 8, respectively, for clarity.
-Compared with the baseline, our judgment agent strategy for correction improves the scores by 6.2% on average when the maximum number of corrections is set to 3.
-Furthermore, it still has 5.3% and 2.6% increases when compared to the other two correction strategies.
-Besides, our strategy achieves the best performance with fewer average number of corrections, indicating high efficiency and accuracy.
-In terms of the impact of the number of generated plans $N_{p}$ and the number of candidate executions $N_{e}$, it is evident that higher values of the two parameters cause superior performance.
-Max Num of Corrections Avg Num of
-3 5 7 10 Corrections
-Baseline 40.2 -
-Internal Correction 40.8 42.0 41.5 38.2 2.8
-Agent for Judgment and Correction 45.1 46.7 44.6 41.2 2.3
-Independent Judgment Agent (Ours) 47.6 47.5 45.3 42.9 1.9
-Baseline 68.8 -
-Internal Correction 69.5 70.3 69.6 64.2 2.3
-Agent for Judgment and Correction 72.2 72.6 71.9 68.9 1.6
-Independent Judgment Agent (Ours) 74.8 74.7 73.7 71.2 1.3
-Table 7: Comparisons among the three correction strategies of various maximum numbers of corrections and the average number of corrections when the limitation is set to 3.
-Line with denotes the results on the three selected benchmarks, while another line is the average results on all benchmarks.
-$N_{p}$ $N_{e}$ 1 2 4 8 16
-1 70.2 70.8 72.0 72.5 72.8
-2 71.1 - - - 73.4
-4 71.5 - - - 73.9
-8 72.0 - - - 74.5
-16 72.2 72.9 73.4 74.1 74.8
-Table 8: Results of different values of $N_{p}$ and $N_{e}$ on all benchmarks.
+Additionally, agent-wise adaptive test-time scaling and mixed reward modeling further reduce the cognitive overload adaptively.
+Extensive comparative experiments and supplementary analysis validate the superiority of our MACT.
+This procedural paradigm constitutes a meaningful exploration of multi-agent frameworks and adaptive test-time scaling strategies for document-based scenarios and tasks.